### v1

# Inference-Time Search using Side Information for Diffusion-based Image Reconstruction

Abstract
Diffusion models have emerged as powerful priors for solving inverse problems. However, existing approaches typically overlook side information that could significantly improve reconstruction quality, especially in severely ill-posed settings. In this work, we propose a novel inference-time search algorithm that guides the sampling process using the side information in a manner that balances exploration and exploitation. This enables more accurate and reliable reconstructions, providing an alternative to the gradient-based guidance that is prone to reward-hacking artifacts. Our approach can be seamlessly integrated into a wide range of existing diffusion-based image reconstruction pipelines. Through extensive experiments on a number of inverse problems, such as box inpainting, super-resolution, and various deblurring tasks including motion, Gaussian, nonlinear, and blind deblurring, we show that our approach consistently improves the qualitative and quantitative performance of diffusion-based image reconstruction algorithms. We also show the superior performance of our approach with respect to other baselines, including reward gradient-based guidance algorithms. The code is available at this repository.
1 Introduction
Diffusion models (Ho et al.,, 2020; Song et al., 2021b, ) have demonstrated remarkable success in generative tasks across various fields like text-to-image synthesis (Rombach et al.,, 2022), protein sequence (Wu et al.,, 2024), video (Ho et al.,, 2022), audio (Kong et al.,, 2021), and language modeling (Austin et al.,, 2021; Sahoo et al.,, 2024). Aside from generation, these models have also shown great promise in solving inverse problems, where the goal is to reconstruct an image from partial or noisy observations (Chung et al., 2023b, ; Song et al., 2023a, ; Rout et al.,, 2023; Song et al.,, 2024; He et al.,, 2024; Ye et al.,, 2024; Zhang et al.,, 2024). Inverse problems differ fundamentally from standard generative tasks (e.g., text-to-image synthesis, personalized editing, style transfer): whereas those tasks are often judged subjectively, inverse problems have a precise objective, to recover a specific ground-truth signal from incomplete measurements. Consequently, fidelity to the ground truth and rigorous quantitative evaluation are critical.
When the observation is heavily degraded, the inverse problem becomes highly ill-posed as many distinct signals can explain the data almost equally well. In this regime, unconstrained posterior sampling rarely recovers the ground truth; it tends instead to produce diverse yet semantically inconsistent reconstructions. A practical solution is to incorporate side information, auxiliary measurements correlated with the target signal, to constrain the solution space and steer the reconstruction toward faithful outcomes. This idea is well established in the classical signal processing literature, where certain structural or encoded properties are used to guide the iterative algorithms that solve the inverse problem (Jones,, 2009; Chun et al.,, 2012; Oymak et al.,, 2013; Ehrhardt et al.,, 2014; Mota et al.,, 2017; Hyder et al.,, 2019). In medical imaging, leveraging complementary measurements or modalities, such as multiple MRI contrasts, multimodal microscopy, or RGB guidance for NIR imaging, has been shown to substantially improve quality (Atalık et al.,, 2025; Tsiligianni and Deligiannis,, 2019).
Figure 1: Illustration of the performance of our inference-time search algorithm for using side information in solving inverse problems, compared with the DPS algorithm (Chung et al., 2023b, ).
While the existing works on diffusion-based solvers have made significant progress on measurement-only inverse problems, they largely sidestep the harder and increasingly common setting where we must also exploit side information (e.g., a reference photograph of the same person, a text description, or features from another modality). A key obstacle is the challenge of learning the conditional distribution $p_{X\mid Y,S}$, where $X$ denotes the target image, $Y$ denotes the noisy measurement, and $S$ denotes the side information. While some recent works (Kim et al., 2025a, ; Chung et al.,, 2025) address the limited setting of textual side information, these approaches typically train a diffusion model to take a specific side-info modality as input; this demands large paired datasets and expensive training, ties the solver to a single conditioning format, and is impractical when the test-time side information differs from what the model was trained on. This motivates us to address the following question:
How can we leverage a pre-trained (unconditional) diffusion prior to solve inverse problems with side information at inference time, without any retraining, so that the method is modality-agnostic and can use text, images, or features depending on the end-use applications? We provide constructive solutions to these questions in our work. Our main contributions are the following.
  • Modeling: We introduce a general modeling approach that incorporates arbitrary side information via an auxiliary reward, characterizing $p_{X\mid S}$ as a reward-tilted version of the pre-trained diffusion prior. This abstraction cleanly decouples the measurement model from the side information, is modality-agnostic (text, image, features), and requires no retraining. We use this modeling with tractable approximations and appropriate error bounds for computing the conditional score functions that are needed for sampling from the pre-trained diffusion models.
  • Algorithm: Motivated by recent successes of inference-time search in LLMs (Snell et al.,, 2025; Setlur et al.,, 2025; Liu et al.,, 2025), we propose a compute-aware, training-free inference-time search framework that can leverage the side information to solve inverse problems. We instantiate this framework by proposing two specific search algorithms: $(i)$ Greedy Search (GS), a strategy that resamples greedily at each step, and $(ii)$ Recursive Fork-Join Search (RFJS), which balances exploration and exploitation through a group-based sampling at each step. The framework operates as a plug-in on top of any standard inverse-problem solvers and supports black-box, non-differentiable rewards. To our knowledge, our work is the first to propose inference-time search with side information for diffusion-based inverse problems.
  • Experiments: We provide extensive experimental evaluations of our proposed approach across linear and nonlinear problems (e.g., box inpainting, super-resolution, motion/Gaussian/nonlinear deblurring) and side-information types (images and text), and demonstrate that our approach outperforms multiple relevant baseline algorithms.
2 Related work
[…]
Inverse problems with side information: Many works in signal processing (Mota et al.,, 2017; Oymak et al.,, 2013; Jones,, 2009; Chun et al.,, 2012; Ehrhardt et al.,, 2014; Hyder et al.,, 2019) integrate structural correlations from auxiliary signals, often via designing appropriate optimization algorithms. In MRI, LeSITA (Tsiligianni and Deligiannis,, 2019) learns coupled sparse representations, and TGVN (Atalık et al.,, 2025) constrains ambiguous subspaces with additional contrasts using learned unrolled networks. Diffusion-based approaches include training with joint priors across modalities (Levac et al.,, 2023; Efimov et al.,, 2025), metadata conditioning (Chung et al.,, 2025), and text-guided regularization (Kim et al., 2025a, ). Most approaches, however, are training-based or bound to one modality of side information associated with the trained conditional diffusion model.
Reward-gradient guidance: LGD (Song et al., 2023b, ) refines DPS via Monte Carlo estimates, while UGD (Bansal et al.,, 2024), FreeDoM (Yu et al.,, 2023), and RB-Modulation (Rout et al.,, 2025) propose to guide the diffusion with a gradient of the reward function. In addition to being gradient-based approaches, they are typically used for semantic generation tasks rather than inverse problems.
SMC methods: Sequential Monte Carlo approaches (Cardoso et al.,, 2024; Dou and Song,, 2024; Wu et al.,, 2023) generate and resample particles under tilted distributions, offering gradient-free alternatives but limited performance at small $N$. DAS (Kim et al., 2025b, ) combines resampling with gradients for text-to-image tasks. These methods rely only on the measurement to guide the unconditional sampler and do not exploit side information.
Inference-time search: Reward-guided inference-time search has advanced LLM reasoning using Process Advantage Verifiers (PAVs) (Setlur et al.,, 2025), compute-optimal scheduling (Snell et al.,, 2025), and reward-guided small models (Liu et al.,, 2025). Some recent works (Singhal et al.,, 2025; Li et al.,, 2025) apply reward-based search in diffusion for text-to-image/protein generation, but do not consider side information or inverse problems.
3 Preliminaries and Problem Formulation
3.1 Preliminaries
Diffusion models: Diffusion models (Ho et al.,, 2020; Song et al., 2021b, ) are powerful generative models that enable sampling from an (unknown) distribution through an iterative process. Diffusion models comprise a forward diffusion process and a reverse denoising process. During the forward process, a clean sample from the distribution $p_{\mathrm{data}}$ is progressively corrupted by the addition of Gaussian noise at each timestep, transforming the data distribution into pure noise. Conversely, the reverse process trains a denoising neural network to iteratively remove this introduced noise, enabling the reconstruction of samples from the initial data distribution. The forward process is represented by the stochastic differential equation (SDE), $\mathrm{d}{\mathbf{x}}_{t}=f({\mathbf{x}}_{t},t)\mathrm{d}t+g(t)\mathrm{d}{\mathbf{w}}_{t},~\forall t\in[0,T],$ where $x_{0}$ is sampled from $p_{\mathrm{data}}$ and $\mathbf{w}_{t}$ is a Wiener process. Common choices for $f,g$ are $f(\mathbf{x}_{t},t)=-({\beta(t)}/{2})\mathbf{x}_{t}$ and $g(t)=\sqrt{\beta(t)}$ for some non-negative monotonic increasing function $\beta(\cdot)$ over $[0,T]$. The corresponding reverse process of this SDE is described by (Anderson,, 1982; Song et al., 2021b, ) $\mathrm{d}\mathbf{x}_{t}=\left(f(\mathbf{x}_{t},t)-g^{2}(t)\nabla_{\mathbf{x}_{t}}\log p_{t}(\mathbf{x}_{t})\right)\mathrm{d}t+\mathrm{d}\mathbf{w}_{t},~\forall t\in[T,0],$ where $p_{t}$ denotes the marginal probability distribution of $\mathbf{x}_{t}$, $\mathbf{x}_{T}$ is sampled according to a standard Gaussian distribution, and $\nabla_{\mathbf{x}_{t}}\log p_{t}(\mathbf{x}_{t})$ represents the score function. Since the marginal distribution $p_{t}$ is unknown, the score function is approximated by a neural network $\mathcal{D}_{\boldsymbol{\theta}}(\mathbf{x}_{t},t)$ via the minimization of a score-matching objective. In practical implementations, the SDE is discretized into $T$ steps, and we define $\alpha_{t}\triangleq\prod_{s=1}^{t}(1-\beta_{s})$.
Solving inverse problems using diffusion models: An inverse problem consists of recovering an unknown signal $\mathbf{x}_{0}$ from noisy, partial observations $\mathbf{y}=\mathbf{A}(\mathbf{x}_{0})+\sigma_{y}\mathbf{z}$, where $\mathbf{A}$ is the measurement model, $\sigma_{y}$ is the observation noise level, and $\mathbf{z}$ is typically a Gaussian noise. Often, $\mathbf{A}$ is non-injective, i.e., multiple signals $\mathbf{x}_{0}$ can produce the same measurement $\mathbf{y}$. A standard approach for estimating $\mathbf{x}_{0}$ is via the Bayesian framework, assuming a prior distribution $p_{0}$ over the signal $\mathbf{x}_{0}$, and sampling from the posterior distribution $\mathbf{x}_{0}\sim p_{0\mid Y}(\cdot\mid\mathbf{y})$. Though $p_{0\mid Y}(\cdot\mid\mathbf{y})$ is not known, this sampling can be achieved by running the backward SDE with replacing the original score function with the conditional score function $\nabla_{\mathbf{x}_{t}}\log p_{t\mid Y}(\mathbf{x}_{t}\mid\mathbf{y})$. Using Bayes’ theorem, $\nabla_{\mathbf{x}_{t}}\log p_{t\mid Y}(\mathbf{x}_{t}\mid\mathbf{y})=\nabla_{\mathbf{x}_{t}}\log p_{t}(\mathbf{x}_{t})+\nabla_{\mathbf{x}_{t}}\log p_{Y\mid t}(\mathbf{y}\mid\mathbf{x}_{t}).$ While the score function network $\mathcal{D}_{\boldsymbol{\theta}}$ of the pre-trained diffusion model can be used to approximate the first term, approximating the second term is significantly more challenging, and numerous approaches (Daras et al.,, 2024) have been proposed to tackle this challenge. In particular, Diffusion Posterior Sampling (DPS) (Chung et al., 2023b, ) proposes a simple approach to approximate $p_{Y\mid t}$ as $p_{Y\mid t}(\mathbf{y}\mid\mathbf{x}_{t})=\mathbb{E}_{\mathbf{x}_{0}\sim p_{0\mid t}(\cdot\mid\mathbf{x}_{t})}[p_{Y\mid 0}(\mathbf{y}\mid\mathbf{x}_{0})]\approx p_{Y\mid 0}(\mathbf{y}\mid\mathbb{E}_{\mathbf{x}_{0}\sim p_{0\mid t}(\cdot\mid\mathbf{x}_{t})}[\mathbf{x}_{0}])$, by pushing the expectation inside the nonlinear $p_{Y\mid 0}(\mathbf{y}\mid\cdot)$. The remaining challenge is to compute the conditional mean $\mathbb{E}_{\mathbf{x}_{0}\sim p_{0\mid t}(\cdot\mid\mathbf{x}_{t})}[\mathbf{x}_{0}]\triangleq\hat{\mathbf{x}}_{0\mid t}(\mathbf{x}_{t})$, which is typically tackled by using Tweedie’s formula (Efron,, 2011), leveraging the fact that $\mathbf{x}_{t}$ given $\mathbf{x}_{0}$ is Gaussian. This results in the estimate
$\hat{\mathbf{x}}_{0\mid t}(\mathbf{x}_{t})=({1}/{\sqrt{\alpha_{t}}})(\mathbf{x}_{t}+(1-\alpha_{t})\nabla_{\mathbf{x}_{t}}\log p_{t}(\mathbf{x}_{t}))\approx({1}/{\sqrt{\alpha_{t}}})(\mathbf{x}_{t}+(1-\alpha_{t})\mathcal{D}_{\boldsymbol{\theta}}(\mathbf{x}_{t},t)).$ (1)
3.2 Problem formulation: solving inverse problems with side information
In many applications, the observation $\mathbf{y}$ alone is insufficient to identify the latent signal $\mathbf{x}_{0}$; auxiliary side information $\mathbf{s}$ (e.g., a reference image, identity/text embedding, or physics-derived features) can dramatically reduce ambiguity. Formally, when side information $\mathbf{s}$ is available, the goal is to sample from the target conditional distribution $p_{0\mid Y,S}(\cdot\mid\mathbf{y},\mathbf{s})$. A seemingly direct route is to train a conditional diffusion model that accepts $\mathbf{s}$ as input, learn the conditional score function $\nabla_{\mathbf{x}_{t}}\log p_{t\mid S}(\mathbf{x}_{t}\mid\mathbf{s})$, and then approximate the full conditional score $\nabla_{\mathbf{x}_{t}}\log p_{t\mid Y,S}(\mathbf{x}_{t}\mid\mathbf{y},\mathbf{s})=\nabla_{\mathbf{x}_{t}}\log p_{t\mid S}(\mathbf{x}_{t}\mid\mathbf{s})+\nabla_{\mathbf{x}_{t}}\log p_{Y\mid t,S}(\mathbf{y}\mid\mathbf{x}_{t},\mathbf{s})$ through a DPS-style method for the second term, to run the backward SDE. However, this training-based approach is often impractical: it demands large paired datasets $(\mathbf{x}_{0},\mathbf{s})$, which are expensive or impossible to curate; it locks the solver to the training modality of $\mathbf{s}$ (a text-conditioned prior cannot natively exploit an image or spectral feature at test time); and general multi-modal conditioning requires prohibitive data and compute. These constraints motivate a training-free alternative that reuses strong unconditional diffusion priors and uses $\mathbf{s}$ only at inference, preserving modality-agnosticism and avoiding costly data collection.
Designing such a training-free method is technically challenging. First, DPS-style derivations rely on tractable likelihoods (e.g., Gaussian $p_{Y\mid 0}$), whereas realistic $p_{S\mid 0}$ are often non-Gaussian implicitly, complicating conditional-score construction. Second, even for measurement-only guidance, computing the conditional score used in the DPS-style algorithms requires back-propagating through the denoiser at every step. Naively extending to side information forces second-order/Hessian terms through the diffusion network. Third, purely gradient-guided diffusion is brittle: it struggles with non-differentiable or black-box rewards, amplifies early-step errors, and can drift off the data manifold. Inference-time search approaches, which have shown remarkable performance improvement in LLMs (Setlur et al.,, 2025; Liu et al.,, 2025; Snell et al.,, 2025) and text-conditioned diffusion models (Singhal et al.,, 2025; Kim et al., 2025b, ), but have not yet been used for solving the inverse problems, offer a promising path to overcome these challenges. In this context, we address the following questions:
$(i)$ Modeling: How can we realize $p_{0\mid Y,S}$ at inference time, without any retraining, by constructing a surrogate objective that is valid across diverse side-information modalities? $(ii)$ Algorithm: How can we design a plug-and-play inference-time search module that is modality-agnostic, compute-aware, and capable of making global corrections (beyond local gradient steps)?
4 Modeling and Algorithm
4.1 Modeling side information using reward function
Given a side-information signal $\mathbf{s}$ corresponding to an unknown $\mathbf{x}_{0}$, and two candiate reconstructions, $\mathbf{x}_{0}^{1}$ and $\mathbf{x}_{0}^{2}$, a principled way to decide which reconstruction is more truthful is to compare the (unknown) conditional probabilities $p_{0\mid S}(\mathbf{x}_{0}^{1}\mid\mathbf{s})$ and $p_{0\mid S}(\mathbf{x}_{0}^{2}\mid\mathbf{s})$. Directly estimating $p_{0\mid S}$ is intractable in our setting: it is typically non-Gaussian, multi-modal, and depends on the data domain and modality of $\mathbf{s}$. We therefore introduce a reward function $r:\mathbb{R}^{d}\times\mathcal{S}\to\mathbb{R}$ that orders reconstructions given $\mathbf{s}$: if $r(\mathbf{x}_{0}^{1},\mathbf{s})>r(\mathbf{x}_{0}^{2},\mathbf{s})$, then $\mathbf{x}_{0}^{1}$ is deemed more compatible with $\mathbf{x}_{0}$ than $\mathbf{x}_{0}^{2}$. This abstraction aligns with many real-world applications (as shown in our experiments): when $\mathbf{s}$ is a text description of the target image $\mathbf{x}_{0}$, we can use a pre-trained text-image model to score text-image alignment. When $\mathbf{s}$ is a reference image of the same entity (e.g., the same person under different poses/lighting), we can use a pre-trained network to score image-image similarity. Such pre-trained rewards are typically available across datasets, and monotone with respect to the intuitive notion of agreement with $\mathbf{x}_{0}$. In this sense, they serve as practically justified surrogates for comparing $p_{0\mid S}$ without requiring an explicit conditional density model.
Our key modeling choice is to use $r(\cdot,\mathbf{s})$ to implicitly characterize $p_{0\mid S}(\cdot\mid\mathbf{s})$ by tilting the unconditional prior $p_{0}$ toward higher-reward regions. Our approach is inspired by the alignment framework used in LLMs (Ouyang et al.,, 2022; Rafailov et al.,, 2023), where the goal is to generate a sample $\mathbf{x}$ that maximizes some reward $r(\mathbf{x})$, while ensuring that the sampling distribution does not deviate too much from the pre-trained distribution $p_{0}$. This is typically formalized as a KL-regularized reward maximization problem, $\max_{p\in\mathcal{P}}\big(\mathbb{E}_{\mathbf{x}\sim p}[r(\mathbf{x})]-\tau D_{\mathrm{KL}}(p\|p_{0})\big)$, where $\tau>0$ offers the trade-off between the deviation from the prior and reward maximization. This optimization problem admits a closed-form solution, $p^{*}(\mathbf{x})\propto p_{0}(\mathbf{x})\exp(r(\mathbf{x})/\tau)$ (Rafailov et al.,, 2023). Based on this intuition, we make the following modeling assumption: the conditional distribution $p_{0\mid S}$ is approximated as,
$p_{0\mid S}(\mathbf{x}_{0}\mid\mathbf{s})\propto p_{0}(\mathbf{x}_{0})\exp\left(\tfrac{r(\mathbf{x}_{0};\mathbf{s})}{\tau}\right),$ (2)
This assumption: $(i)$ preserves the powerful unconditional diffusion prior $p_{0}$, $(ii)$ injects modality-agnostic side information via a reward, and $(iii)$ produces a tractable objective that we can combine with the measurement model to target $p_{0\mid Y,S}$ at inference time using a pre-trained diffusion model. We do not claim optimality of Eq.˜2; rather, we show it leads to a practical, training-free algorithm that consistently improves reconstructions over strong baselines while keeping compute comparable.
We now leverage Eq.˜2 to compute the conditional posteriors for the reverse diffusion.
Proposition 1.
Let $p_{t\mid t+1,Y,S}$ denote the conditional posterior distribution for the reverse diffusion process. Then using (2) we have
$p_{t\mid t+1,Y,S}(\mathbf{x}_{t}\mid\mathbf{x}_{t+1},\mathbf{y},\mathbf{s})$ $\propto p_{t\mid t+1,Y}(\mathbf{x}_{t}\mid\mathbf{x}_{t+1},\mathbf{y})\exp(V_{t}^{\tau}(\mathbf{x}_{t};\mathbf{s},\mathbf{y})),$ (3)
$p_{t\mid Y,S}(\mathbf{x}_{t}\mid\mathbf{y},\mathbf{s})$ $\propto p_{t\mid Y}(\mathbf{x}_{t}\mid\mathbf{y})\exp(V_{t}^{\tau}(\mathbf{x}_{t};\mathbf{s},\mathbf{y})),$ (4)
where $V_{t}^{\tau}(\mathbf{x}_{t};\mathbf{s},\mathbf{y})\triangleq\log\mathbb{E}_{\mathbf{x}_{0}\sim p_{0\mid t,Y}(\cdot\mid\mathbf{x}_{t},\mathbf{y})}[\exp(r(\mathbf{x}_{0};\mathbf{s})/\tau)]$.
The proof is provided in Appendix A.1. Using (4), we can get the conditional score function as,
$\begin{split}\nabla_{\mathbf{x}_{t}}\log p_{t\mid Y,S}(\mathbf{x}_{t}\mid\mathbf{y},\mathbf{s})&=\nabla_{\mathbf{x}_{t}}\log p_{t}(\mathbf{x}_{t})+\nabla_{\mathbf{x}_{t}}\log p_{Y\mid t}(\mathbf{y}\mid\mathbf{x}_{t})+\nabla_{\mathbf{x}_{t}}V_{t}^{\tau}(\mathbf{x}_{t};\mathbf{s},\mathbf{y}).\end{split}$ (5)
The computation of $V_{t}^{\tau}$ is not straightforward. So, we use a DPS-style approximation as $V_{t}^{\tau}(\mathbf{x}_{t};\mathbf{s},\mathbf{y})=\log\mathbb{E}_{\mathbf{x}_{0}\sim p_{0\mid t,Y}(\cdot\mid\mathbf{x}_{t},\mathbf{y})}[\exp(r(\mathbf{x}_{0};\mathbf{s})/\tau)]\approx r(\mathbb{E}_{\mathbf{x}_{0}\sim p_{0\mid t,Y}(\cdot\mid\mathbf{x}_{t},\mathbf{y})}[\mathbf{x}_{0}];\mathbf{s})/\tau=r(\hat{\mathbf{x}}_{0\mid t,Y}(\mathbf{x}_{t},\mathbf{y});\mathbf{s})/\tau$. Using some approximation and the fact that $p_{Y\mid 0}$ is Gaussian, we can get
$\hat{\mathbf{x}}_{0\mid t,Y}(\mathbf{x}_{t},\mathbf{y})\approx\hat{\mathbf{x}}_{0\mid t}(\mathbf{x}_{t})-({(1-\alpha_{t})}/{\sqrt{\alpha_{t}}})\eta\nabla_{\mathbf{x}_{t}}\lVert\mathbf{y}-\mathbf{A}\hat{\mathbf{x}}_{0\mid t}(\mathbf{x}_{t})\rVert_{2}^{2},$ (6)
$V_{t}^{\tau}(\mathbf{x}_{t};\mathbf{s},\mathbf{y})\approx\hat{V}_{t}^{\tau}(\mathbf{x}_{t};\mathbf{s},\mathbf{y})\triangleq r\left(\hat{\mathbf{x}}_{0\mid t,Y}(\mathbf{x}_{t},\mathbf{y});\mathbf{s}\right)/\tau.$ (7)
In Appendix A.2, we have provided the details of the steps leading to Eq.˜6-Eq.˜7.
We characterize the error in approximating the value function, $|V_{t}^{\tau}(\mathbf{x}_{t};\mathbf{s},\mathbf{y})-\hat{V}_{t}^{\tau}(\mathbf{x}_{t};\mathbf{s},\mathbf{y})|$, in LABEL:{prop:approx-error-bound}, which is deferred to Appendix A.2.
We can now get $\nabla_{\mathbf{x}_{t}}\log p_{t\mid Y,S}(\mathbf{x}_{t}\mid\mathbf{y},\mathbf{s})$ given in Eq.˜5 by replacing $\nabla_{\mathbf{x}_{t}}V_{t}^{\tau}(\mathbf{x}_{t};\mathbf{s},\mathbf{y})$ with $\nabla_{\mathbf{x}_{t}}\hat{V}^{\tau}_{t}(\mathbf{x}_{t};\mathbf{y},\mathbf{s})$. However, running a backward diffusion using $\nabla_{\mathbf{x}_{t}}\hat{V}_{t}^{\tau}(\mathbf{x}_{t};\mathbf{s},\mathbf{y})$ is computationally infeasible because it involves computing second-order derivatives through the denoiser network. This issue, however, can be circumvented by making a further approximation, by setting $\eta=0$ in Eq.˜6 to get $\hat{\mathbf{x}}_{0\mid t,Y}(\mathbf{x}_{t},\mathbf{y})\approx\hat{\mathbf{x}}_{0\mid t}(\mathbf{x}_{t})$, which leads to the approximation $\nabla_{\mathbf{x}_{t}}V_{t}^{\tau}(\mathbf{x}_{t};\mathbf{s},\mathbf{y})\approx\nabla_{\mathbf{x}_{t}}r(\hat{\mathbf{x}}_{0\mid t}(\mathbf{x}_{t});\mathbf{s})$. We show that approximation error remains small when $t$ is small in Appendix A.2 even when $\eta=0$. This approach then reduces to the reward gradient guidance (RGG) approach used for the inference-time alignment of diffusion models (Bansal et al.,, 2024; Kim et al., 2025b, ; Yu et al.,, 2023; He et al.,, 2024), with the critical difference being that the guidance is from both $\mathbf{s}$ and $\mathbf{y}$.
The RGG approach, however, is limited only to differentiable rewards, and even when they are differentiable, calculating a gradient through the denoiser network at each step of the backward diffusion is computationally intensive and can be ill-suited for many end-use edge-device applications. Moreover, the hyperparameter that determines the weight of the reward gradient guidance is highly sensitive and is difficult to tune, leading to limited performance improvements and undesirable artifacts in the reconstructed images. We later illustrate these issues in Section˜B.5. This motivates us to pursue a gradient-free approach for leveraging the side information for inverse problems.
4.2 Inference-time Search Algorithms for Inverse Problems
Inference-time search methods have recently gained traction as a means of improving the quality of output from LLMs (Snell et al.,, 2025; Setlur et al.,, 2025; Liu et al.,, 2025). The key objective of a search algorithm is to solve a multi-step decision-making problem with balanced exploration and exploitation. While Monte Carlo Tree Search (MCTS) (Kocsis and Szepesvári,, 2006) was successful in large-scale reinforcement learning systems like AlphaGo (Silver et al.,, 2016), they are infeasible for diffusion models: estimating the expected reward of a noisy state $\mathbf{x}_{t}$ would require repeated rollouts of the reverse process. Training a value function to guide the search rewards is another alternative, but this demands additional training customized to each modality and type of side information. These limitations motivate inference-time methods that are both training-free and computationally tractable.
Particle-based procedures offer one such approach, using the distribution given by Eq. (3), where the value function is replaced by the approximation in Eq. (7). At a given step, suppose we have $N$ samples $\mathbf{x}_{t+1}[1],\dots,\mathbf{x}_{t+1}[N]\sim p_{t+1\mid Y,S}$. One way to generate samples from $p_{t\mid Y,S}$ is to $(i)$ propose candidates $\tilde{\mathbf{x}}_{t}[i]\sim p_{t\mid t+1,Y}(\cdot\mid\mathbf{x}_{t+1}[i],\mathbf{y})$, $(ii)$ compute rewards $r[i]=r(\hat{\mathbf{x}}_{0\mid t,Y}(\tilde{\mathbf{x}}_{t}[i],\mathbf{y});\mathbf{s})$ (approximate value) $(iii)$ assign weights $w[i]\propto\exp(r[i]/\tau)$ and resample indices with replacement $I[i]\sim\mathrm{Cat}(w[1:N])$, and $(iv)$ retain $\mathbf{x}_{t}[I[i]]$ for the next step. In theory, such particle methods converge to the target distribution as $N\to\infty$ and with exact tilting (Wu et al.,, 2023; Dou and Song,, 2024). In practice, however, finite $N$ and approximate tilting has some issues: frequent resampling favors exploitation but risks reward over-optimization, while no resampling preserves data consistency but requires prohibitively many particles to harness and optimize the reward.
To address this trade-off, we modify the step $(iii)$ by introducing grouped resampling at each time step $t$, where particles are resampled within groups of size $g_{t}$. Formally, let the index set of the $i$-th group be $\mathcal{G}_{i}=\{(i-1)g_{t}+1,\dots,ig_{t}\}$ for $i=1,\dots,N/g_{t}$. For each group, we draw $g_{t}$ indices with replacement according to the weights within that groups, i.e., $I^{\prime}[j]\sim\mathrm{Cat}(w[\mathcal{G}_{i}])\in[1:g_{t}]$. The absolute indices are then obtained by shifting $I^{\prime}$ as $I[(i-1)g_{t}+j:ig_{t}+j]=(i-1)g_{t}+I^{\prime}[j]$ for $j=1,\dots,g_{t}$. Based on the choice of $g_{t}$, we introduce two specific search strategies.
Greedy Search (GS): Here, we use a fixed resampling period $B$ and select $g_{t}=N$ whenever $t\bmod B=0$, and $g_{t}=1$ otherwise. Greedy Search reduces to the Best-of-N (BON) strategy when $B\geq T$, since in that case $g_{t}=1$ for all $t$. Smaller values of $B$ emphasize short-term reward exploitation, while larger values promote long-term consistency and exploration. An illustration of Greedy Search, with resampling interval $B$, is provided in Figure 2, where the particles evolve independently between resampling events and only interact at steps that are multiples of $B$.
Recursive Fork-Join Search (RFJS): Greedy search considers the largest resampling group size ($N$) at fixed time periods of $B$, and greedily selects one particle from this group, which leads to an exploitation-style approach in search. Selecting the smallest group size ($g_{t}=1$) leads to a pure exploration-style search of BON. Ideally, one should combine the benefits of resampling with multiple group sizes at multiple time steps to get a balanced exploration and exploitation.
Figure 2: Illustration of the group size resampling strategies of different search algorithms.
To this end, we propose a recursive grouping and sampling approach, which we call recursive fork-join search (RFJS), in which the resampling group sizes vary systematically over time. At every $B$ steps, all $N$ particles are resampled together; at every $B/2$ steps, the particles are partitioned into groups of size $N/2$ that are resampled independently; at every $B/4$ steps, groups of size $N/4$ are resampled; and so on. This hierarchical schedule is illustrated in Figure 2. As a concrete example, consider $N=8$. In this case, groups of size at least $N/4=2$ are resampled every $B/4$ steps, groups of size $N/2=4$ are resampled every $B/2$ steps, and all $N=8$ particles are resampled every $B$ steps. When multiple group sizes are scheduled to be resampled at the same time step $t$, the larger group size always takes precedence. For example, although $t=B/2$ is also a multiple of $B/4$, the scheme prioritizes the larger group size. Thus, rather than resampling groups of size $2$, we resample groups of size $4$ at $t=B/2$. Similarly, at $t=B$, the entire set of $N=8$ particles is resampled jointly. More generally, the group size at time step $t$ is given by $g_{t}=N/2^{j^{*}}$, where $j^{*}=\min\{\,i\geq 0:\;t\bmod(B/2^{i})=0\,\}$. The localized resampling (fork) at intermediate group sizes encourages balanced exploration, while the recursive return to larger group sizes (join) encourages exploitation. Naively reducing $B$ in GS does not balance this trade-off well and may lead to an undesirable compromise between exploration and exploitation.
We have summarized this inference-time search framework in Algorithm 1 in the Appendix A.3. Our framework is modular: the resampling rule, whether BON, GS, or RFJS, can be chosen depending on budget and application. Since this requires no retraining and works with arbitrary reward functions, it can be incorporated into any diffusion-based inverse problem solvers with minimal modification.
5 Experiments
Side Measurement DPS RFJSearch (Ours) Ground Truth Side Measurement DPS RFJSearch (Ours) Ground Truth Box Inpainting Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Super Resolution Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Motion Deblur Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Gaussian Deblur Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Nonlinear Deblur Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Blind Deblur Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Figure 3: Image as side information: Qualitative illustration of the performance of our RFJS algorithm compared to the DPS baseline on linear and nonlinear inverse problems. RFJS is able to capture many details that are missed by the DPS baseline to achieve a superior reconstruction quality.
 Box Inpainting  Super Resolution ($\times$4)  Non-linear Deblur
Algorithm FS ($\downarrow$) PSNR ($\uparrow$) LPIPS ($\downarrow$) SSIM ($\uparrow$) FS ($\downarrow$) PSNR ($\uparrow$) LPIPS ($\downarrow$) SSIM ($\uparrow$) FS ($\downarrow$) PSNR ($\uparrow$) LPIPS ($\downarrow$) SSIM ($\uparrow$)
RFJS (ours) $\mathbf{0.308}$ $\mathbf{28.29}$ $\mathbf{0.136}$ $\mathbf{0.855}$ $\mathbf{0.380}$ $\mathbf{25.26}$ $\mathbf{0.225}$ $\underline{0.695}$ $\mathbf{0.394}$ $\underline{23.89}$ $\mathbf{0.229}$ $\underline{0.668}$
GS (ours) $\underline{0.349}$ $\underline{28.22}$ $\underline{0.137}$ $\underline{0.855}$ $\underline{0.460}$ $\underline{25.24}$ $\underline{0.225}$ $\mathbf{0.696}$ $\underline{0.467}$ $\mathbf{23.92}$ $0.232$ $\mathbf{0.669}$
RGG $0.475$ $27.96$ $0.138$ $0.851$ $0.573$ $25.13$ $0.228$ $0.690$ $0.654$ $23.89$ $\underline{0.231}$ $0.665$
BON $0.584$ $28.20$ $0.137$ $0.854$ $0.915$ $25.14$ $0.229$ $0.694$ $0.881$ $23.89$ $0.233$ $0.667$
DPS $0.739$ $27.93$ $0.139$ $0.852$ $1.042$ $25.13$ $0.229$ $0.693$ $1.008$ $23.87$ $0.232$ $\underline{0.666}$
 Motion Deblur Gaussian Deblur Blind Deblur
RFJS (ours) $\mathbf{0.326}$ $\mathbf{26.64}$ $\mathbf{0.193}$ $\mathbf{0.736}$ $\mathbf{0.330}$ $\mathbf{26.20}$ $\mathbf{0.196}$ $\mathbf{0.712}$ $\mathbf{0.341}$ $\underline{25.04}$ $\mathbf{0.209}$ $\underline{0.707}$
GS (ours) $\underline{0.392}$ $\underline{26.58}$ $\underline{0.193}$ $\underline{0.735}$ $\underline{0.385}$ $\underline{26.16}$ $\underline{0.198}$ $\underline{0.711}$ $\underline{0.417}$ $25.04$ $0.211$ $0.706$
RGG $0.497$ $26.55$ $0.193$ $0.733$ $0.495$ $26.15$ $0.200$ $0.709$ $0.473$ $24.97$ $0.211$ $0.701$
BON $0.671$ $26.57$ $0.194$ $0.735$ $0.667$ $26.18$ $0.201$ $0.711$ $0.642$ $\mathbf{25.15}$ $\underline{0.210}$ $\mathbf{0.708}$
DPS $0.815$ $26.54$ $0.194$ $0.734$ $0.807$ $26.15$ $0.200$ $0.711$ $0.779$ $24.98$ $0.213$ $0.704$
Table 1: Image as side information: Quantitative comparison of our GS and RFJS algorithms with baseline algorithms. For each evaluation metric, the best result is shown in bold, and the second best is underlined. RFJS and GS achieve superior performance consistently across all tasks and metrics.
Measurement DPS RFJSearch (Ours) Ground Truth Measurement DPS RFJSearch (Ours) Ground Truth Box Inpainting Refer to caption Refer to caption Refer to caption Refer to caption Super Resolution Refer to caption Refer to caption Refer to caption Refer to caption Motion Deblur Refer to caption Refer to caption Refer to caption Refer to caption Gaussian Deblur Refer to caption Refer to caption Refer to caption Refer to caption Nonlinear Deblur Refer to caption Refer to caption Refer to caption Refer to caption Blind Deblur Refer to caption Refer to caption Refer to caption Refer to caption
Figure 4: Text as side information : Qualitative illustration of the performance of our RFJS algorithm compared to the DPS baseline. For example, the side information provided for the super resolution task is ‘golden retriever sitting on a snowy frozen lake, facing forward’. RFJS is able to capture many details that are missed by the DPS baseline to achieve a superior reconstruction quality.
Box Inpainting Super Resolution ($\times$32) Non-linear Deblur
Algorithm CS ($\uparrow$) PSNR ($\uparrow$) SSIM ($\uparrow$) LPIPS ($\downarrow$) CS ($\uparrow$) PSNR ($\uparrow$) SSIM ($\uparrow$) LPIPS ($\downarrow$) CS ($\uparrow$) PSNR ($\uparrow$) SSIM ($\uparrow$) LPIPS ($\downarrow$)
RFJS (ours) $\mathbf{0.901}$ $\mathbf{20.75}$ $\mathbf{0.678}$ $\mathbf{0.294}$ $\mathbf{0.801}$ $17.13$ $\mathbf{0.352}$ $\mathbf{0.4926}$ $\underline{0.863}$ $\mathbf{20.58}$ $\mathbf{0.473}$ $\mathbf{0.405}$
GS (ours) $\underline{0.894}$ $19.76$ $\underline{0.676}$ $\underline{0.305}$ $\underline{0.791}$ $\underline{17.20}$ $\underline{0.351}$ $0.5094$ $\mathbf{0.865}$ $20.32$ $0.456$ $0.405$
BON $0.882$ $\underline{19.99}$ $0.672$ $0.308$ $0.788$ $\mathbf{17.21}$ $0.350$ $\underline{0.5003}$ $0.855$ $20.52$ $0.464$ $0.406$
DPS $0.871$ $19.86$ $0.672$ $0.312$ $0.731$ $16.90$ $0.330$ $0.5220$ $0.839$ $20.55$ $0.469$ $0.409$
Motion Deblur Gaussian Deblur Blind Deblur
RFJS (ours) $\mathbf{0.858}$ $\underline{18.61}$ $\underline{0.402}$ $\mathbf{0.424}$ $\mathbf{0.843}$ $\mathbf{18.10}$ $\underline{0.358}$ $\underline{0.457}$ $\mathbf{0.851}$ $\underline{18.84}$ $\underline{0.412}$ $\mathbf{0.433}$
GS (ours) $0.835$ $17.83$ $0.369$ $0.453$ $\underline{0.835}$ $17.96$ $0.356$ $0.457$ $\underline{0.835}$ $\mathbf{18.93}$ $\mathbf{0.414}$ $\underline{0.438}$
BON $\underline{0.848}$ $\mathbf{19.24}$ $\mathbf{0.415}$ $\underline{0.427}$ $0.831$ $\underline{17.99}$ $\mathbf{0.365}$ $\mathbf{0.452}$ $0.831$ $18.78$ $0.410$ $0.443$
DPS $0.794$ $18.16$ $0.384$ $0.458$ $0.778$ $16.79$ $0.329$ $0.487$ $0.793$ $18.82$ $0.409$ $0.459$
Table 2: Text as side information: Quantitative comparison of our GS and RFJS algorithms with baseline algorithms. RFJS and GS achieve better performance across all tasks and metrics.
Side Measurement ContextMRI RFJSearch (Ours) Ground Truth PDFS with PD Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption PD with PDFS Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Figure 5: Contrast Image as Side Information: Qualitative MRI reconstruction with RFJS vs. ContextMRI. The shapes and line edges are well preserved in our reconstruction.
 PDFS with PD
Algorithm PSNR ($\uparrow$) SSIM ($\uparrow$) LPIPS ($\downarrow$) NMI ($\uparrow$)
RFJS 25.85 0.801 0.375 0.457
GS 25.33 0.797 0.375 0.455
BON 25.47 0.797 0.376 0.454
ContextMRI 25.39 0.795 0.383 0.451
 PD with PDFS
RFJS 27.85 0.920 0.358 0.579
GS 27.80 0.920 0.360 0.579
BON 27.80 0.918 0.366 0.570
ContextMRI 27.46 0.915 0.375 0.563
Figure 6: Quantitative MRI reconstruction results (fastMRI knee, AF=16, ACS=2%).
5.1 Experimental Setup
We evaluate our inference-time search framework for solving inverse problems with side information by instantiating two specific search algorithms we proposed: Greedy Search (GS) and Recursive Fork Join Search (RFJS), both of which are described in the previous section. We consider two types of side information: $(i)$ image as side information, where a reference image of the same entity (here, the same person under different poses/lighting) is used as side information, and $(ii)$ text as side information, where a text description of the target image is used as a side information.
We demonstrate the plug-and-play nature of our algorithms by considering four different baseline inverse problem solvers: $(i)$ DPS (Chung et al., 2023b, ), $(ii)$ BlindDPS (Chung et al., 2023a, ), $(iii)$ MPGD (He et al.,, 2024), and $(iv)$ DAPS (Zhang et al.,, 2024). Due to page limitation, the evaluation results using DAPS and MPGD are deferred to Appendix B.
Inverse problems: We evaluate our algorithms on six inverse problems, covering both linear and nonlinear problems. The linear problems are: $(i)$ box inpainting, $(ii)$ super resolution, $(iii)$ motion deblurring, and $(iv)$ Gaussian deblurring. The nonlinear problems are: $(v)$ nonlinear deblurring, and $(vi)$ blind deblurring. A detailed description of these inverse problems is given in Appendix C.
Baselines: We compare the performance of GS and RFJS against the following baselines: $(i)$ Baseline solvers (DPS, BlindDPS, MPGD, DAPS), $(ii)$ Best-of-N (BoN), which generates $N$ independent samples and selects the one with the best reward at the end, $(iii)$ Reward Gradient Guidance (RGG), which solves the inverse problem by running the backward diffusion according to Eq.˜5, but with the approximation $\nabla_{\mathbf{x}_{t}}V_{t}^{\tau}(\mathbf{x}_{t};\mathbf{s},\mathbf{y})\approx\nabla_{\mathbf{x}_{t}}r(\hat{\mathbf{x}}_{0\mid t}(\mathbf{x}_{t});\mathbf{s})$. Unless otherwise noted, hyperparameters, including guidance scale, number of diffusion steps, and task-specific settings, match the original baseline implementations. The specific values of hyperparamaters are listed in Appendix C. All the experiments are run on NVIDIA A100 GPUs on an internal compute cluster.
5.2 Main Results
Refer to caption Refer to caption Refer to caption    Refer to caption Refer to caption Refer to caption
Ground truth Input RFJS (Ours) DPS (1) DPS (2) DPS (3)
Figure 7: Qualitative illustration of the relevance of the FaceSimilarity metric and the superior performance of RFJS in identity preservation. RFJS reconstruction is clearly more faithful to the ground truth, yet PSNR, SSIM, and LPIPS values slightly favor the BlindDPS outputs.
Image as side information: The goal is to reconstruct a human face from a noisy observation when another image of the same identity is available (Fig. 3). Using Celeb-HQ (Na et al.,, 2022) as an out-of-distribution set and a diffusion model pretrained on FFHQ (Chung et al., 2023b, ), we sample two random images per identity for target and side information. We compute the reward as follows: first, detect the face using MTCNN (Zhang et al.,, 2016) and then extract identity features with AdaFace (Kim et al.,, 2022). Then, we measure the reward as the negative of the FaceSimilarity (FS) loss, computed as the distance between the identity embeddings of the reconstructed and side-information faces, extracted by pretrained AdaFace network. We evaluate on 64 pairs, using $N=8$ particles and $B=16$, with a gradient scale $0.5$ for RGG. We evaluate with standard metrics, PSNR, SSIM, and LPIPS, but these often fail to measure the identity similarity. Thus, we use FaceSimilarity (FS), comparing the reconstruction to the ground truth for a more reliable measure of identity preservation. Table 1 shows that both proposed inference-time search methods, GS and RFJS, outperform baselines, with RFJS achieving the best overall scores indicating a stronger balance between exploration and exploitation. Qualitative results given in Fig 3 show sharper facial details and preserved identity traits indicating the importance of the FS metric. Fig 7 provides a detailed qualitative example highlighting why FaceSimilarity is essential in addition to standard metrics such as PSNR, SSIM, and LPIPS. It compares three samples generated by the BlindDPS algorithm with our RFJS reconstruction. Although our result better preserves identity and achieves a higher FaceSimilarity score, its PSNR and LPIPS are worse than BlindDPS, underscoring that traditional metrics may not capture semantic fidelity. This illustrates the need for task-specific metrics in inverse problems to ensure qualitative performance.
Text as side information: The goal is to reconstruct an image from its noisy observation, with a text description of the image available as side information. We use a pre-trained diffusion model trained on the ImageNet data (Dhariwal and Nichol,, 2021). We use 25 images from the ImageNet validation set to evaluate the algorithms and generated a short one-sentence textual description for each image using ChatGPT. We use ImageReward (Xu et al.,, 2023), a pre-trained network that measures text-to-image similarity, as the reward function. We consider some inverse problem tasks that are significantly challenging, including $\times 32$ super resolution, and strong blur with larger kernels. Experiments use N=4N{=}4 and B=100B{=}100, and we report the standard metrics and CLIPScore (Radford et al.,, 2021). CLIPScore measures the cosine similarity between CLIP image embeddings of the ground truth and reconstruction, providing a semantically informed metric that reflects both visual and textual alignment. It can be seen in Fig. 4 that the qualitative reconstructions closely match textual descriptions. The quantitative metrics are in Table 2 where both GS and RFJS outperform competing baselines, with RFJS achieving the highest CLIPScore.
MRI with multi-contrast side information: Finally, we test on fastMRI knee dataset (Zbontar et al.,, 2018) with the ContextMRI model (Chung et al.,, 2025). We pair PD and PDFS contrasts, reconstructing one from the other under highly accelerated $16\times$ undersampling with 2% ACS. We use normalized mutual information (NMI) as reward, which is robust to contrast changes. Table 6 shows our methods consistently outperform the baseline in all the metrics of interest. Figure 6 highlights sharper edges and more faithful structure.
Additional Experiments: To demonstrate the generality of our framework, we extended our search algorithm beyond DPS to other samplers, including DAPS and MPGD. Qualitative and quantitative results for these experiments are provided in Appendix B, along with additional DPS results for both types of side information. To further build intuition and to compare the performance of RFJ and Greedy Search, Appendix B.4 presents 2D toy examples illustrating the benefits of side information and the impact of $B$; Fig. 12 shows that RFJS is more robust than GS when the reward is non-linear and non-convex. We also conducted several ablations to analyze the role of side information and the scalability of our approach. Appendix B.5 studies the sensitivity of the gradient-guided methods. Appendix B.6 examines the effect of the number of particles: increasing $N$ improves exploration and reward, while runtime grows sublinearly due to parallelization (Appendix B.7). Hyperparameter details are summarized in Appendix C.
6 Conclusion
We proposed a lightweight, modular inference-time search algorithm that integrates side information into diffusion-based image reconstruction, in a principled way. By adaptively guiding the generative process, our method delivers substantial quality gains, especially in ill-posed settings, while requiring only minimal changes to existing pipelines. Extensive experiments across standard reconstruction tasks show consistent improvements in both visual fidelity and quantitative metrics, and our approach surpasses gradient-based alternatives. These results highlight the power of leveraging side information at inference time to make diffusion-based solvers more reliable and accurate.
Appendices
Appendix A Proofs
A.1 Proof of Proposition 1
To begin, recall that $p_{0\mid S}(\mathbf{x}_{0}|\mathbf{s})=\frac{1}{Z}p_{0}(\mathbf{x}_{0})e^{\frac{r(\mathbf{x}_{0};\mathbf{s})}{\tau}}$. Using Bayes’ rule, we can rewrite this expression as
$\frac{1}{Z}p_{0}(\mathbf{x}_{0})\exp\left(\frac{r(\mathbf{x}_{0};\mathbf{s})}{\tau}\right)$ $=\frac{p_{0,S}(\mathbf{x}_{0},\mathbf{s})}{p_{S}(\mathbf{s})}=\frac{p_{S\mid 0}(\mathbf{s}\mid\mathbf{x}_{0})p_{0}(\mathbf{x}_{0})}{{p_{S}(\mathbf{s})}}$
and we gather that, for $\mathbf{s}$ fixed, $p_{S\mid 0}(\mathbf{s}|\mathbf{x}_{0})\propto e^{\frac{r(\mathbf{x}_{0};\mathbf{s})}{\tau}}$ where the proportionality is up to constants on independent of $\mathbf{x}_{0}$. Starting from the LHS of (3), we first apply Bayes’ rule, reverse the conditioning, and introduce a marginalized $\mathbf{x}_{0}$ to leverage conditional independence. This sequence leads to
$p_{t\mid t+1,Y,S}(\mathbf{x}_{t}\mid\mathbf{x}_{t+1},\mathbf{y},\mathbf{s})=\frac{p_{t,S\mid t+1,Y}(\mathbf{x}_{t},\mathbf{s}\mid\mathbf{x}_{t+1},\mathbf{y})}{p_{S\mid t+1,Y}(\mathbf{s}\mid\mathbf{x}_{t+1},\mathbf{y})}$
$=\frac{p_{t\mid t+1,Y}(\mathbf{x}_{t}\mid\mathbf{x}_{t+1},\mathbf{y})p_{S\mid t,t+1,Y}(\mathbf{s}\mid\mathbf{x}_{t},\mathbf{x}_{t+1},\mathbf{y})}{p_{S\mid t+1,Y}(\mathbf{s}\mid\mathbf{x}_{t+1},\mathbf{y})}$
$\propto p_{t\mid t+1,Y}(\mathbf{x}_{t}\mid\mathbf{x}_{t+1},\mathbf{y})\int_{\mathbf{x}_{0}}p_{S\mid t,t+1,Y}(\mathbf{s}\mid\mathbf{x}_{0},\mathbf{x}_{t},\mathbf{x}_{t+1},\mathbf{y})p_{0\mid t,t+1,Y}(\mathbf{x}_{0}\mid\mathbf{x}_{t},\mathbf{x}_{t+1},\mathbf{y})d\mathbf{x}_{0}$
$=p_{t\mid t+1,Y}(\mathbf{x}_{t}\mid\mathbf{x}_{t+1},\mathbf{y})\int_{\mathbf{x}_{0}}p_{S\mid 0,t,t+1,Y}(\mathbf{s}\mid\mathbf{x}_{0},\mathbf{x}_{t},\mathbf{x}_{t+1},\mathbf{y})p_{0\mid t,Y}(\mathbf{x}_{0}\mid\mathbf{x}_{t},\mathbf{y})d\mathbf{x}_{0}$
$\propto p_{t\mid t+1,Y}(\mathbf{x}_{t}\mid\mathbf{x}_{t+1},\mathbf{y})\int_{\mathbf{x}_{0}}p_{S\mid 0}(\mathbf{s}\mid\mathbf{x}_{0})p_{0\mid t,Y}(\mathbf{x}_{0}\mid\mathbf{x}_{t},\mathbf{y})d\mathbf{x}_{0}$
$=p_{t\mid t+1,Y}(\mathbf{x}_{t}\mid\mathbf{x}_{t+1},\mathbf{y})\mathbb{E}_{\mathbf{x}_{0}\sim p_{0\mid t,Y}(\cdot\mid\mathbf{x}_{t},\mathbf{y})}[\exp(r(\mathbf{x}_{0};\mathbf{s})/\tau)]$
$\propto p_{t\mid t+1,Y}(\mathbf{x}_{t}\mid\mathbf{x}_{t+1},\mathbf{y})\exp(V_{t}^{\tau}(\mathbf{x}_{t};\mathbf{s},\mathbf{y})).$
The penultimate step follows from the discussion at the onset of the proof. The last step captures the definition for $V_{t}^{\tau}(\mathbf{x}_{t};\mathbf{s},\mathbf{y})$ found in Proposition 1. The proof of (4) is similar without conditioning on $t+1$.
Proof of Value-titled KL. Given a distribution $p_{0}$ over $\mathbb{R}^{d}$, a reward function $r:\mathbb{R}^{d}\to\mathbb{R}$, and $\tau>0$, we are interested in sampling from the distribution $p^{*}$ given by
$p^{*}$ $=\arg\max_{p}\mathbb{E}_{\mathbf{x}\sim p}[r(\mathbf{x})]-\tau D_{\rm KL}(p\|p_{0})$
$=\arg\max_{p}\mathbb{E}_{\mathbf{x}\sim p}\left[r(\mathbf{x})-\tau\log\frac{p(\mathbf{x})}{p_{0}(\mathbf{x})}\right]$
$=\arg\min_{p}\mathbb{E}_{\mathbf{x}\sim p}\left[\log\frac{p(\mathbf{x})}{p_{0}(x)}-\frac{r(\mathbf{x})}{\tau}\right]$
$=\arg\min_{p}\mathbb{E}_{x\sim p}\left[\log\frac{p(\mathbf{x})}{p_{0}(\mathbf{x})e^{r(\mathbf{x})/\tau}}\right]\triangleq\arg\min_{p}\mathcal{L}(p).$
Let $q(\mathbf{x})\triangleq\frac{1}{Z}p_{0}(\mathbf{x})e^{r(\mathbf{x})/\tau}$, where $Z$ is chosen such that $\int q(\mathbf{x})d\mathbf{x}=1$. Then
$\mathcal{L}(p)=\mathbb{E}_{\mathbf{x}\sim p}\left[\log\frac{p(\mathbf{x})}{p_{0}(\mathbf{x})e^{r(\mathbf{x})/\tau}}\right]=\mathbb{E}_{\mathbf{x}\sim p}\left[\log\frac{p(\mathbf{x})}{Zq(\mathbf{x})}\right]=D_{\rm KL}(p\|q)-\log Z.$
By the non-negativity of KL-divergence, $\mathcal{L}(p)\geq\mathcal{L}(q)$ for any distribution $p$, and so $p^{*}=q$, or $p(\mathbf{x})\propto p_{0}(\mathbf{x})e^{r(\mathbf{x})/\tau}$.
A.2 Value Approximation Bound
In the following, we provide the steps that lead to Eq. 7, and subsequently bound the approximation error. We begin with the following lemma.
Lemma 2.
The conditional mean of $X_{0}$ given $X_{t}=\mathbf{x}_{t}$ and $Y=\mathbf{y}$ is given by
$\hat{\mathbf{x}}_{0\mid t,Y}(\mathbf{x}_{t},\mathbf{y})=\hat{\mathbf{x}}_{0\mid t}(\mathbf{x}_{t})+\left(\frac{1-\alpha_{t}}{\sqrt{\alpha_{t}}}\right)\nabla_{\mathbf{x}_{t}}\log p_{Y\mid t}(\mathbf{y}\mid\mathbf{x}_{t}).$ (8)
Proof. For any distribution over $X_{0}$ since $p_{t\mid 0}(\mathbf{x}_{t}\mid\mathbf{x}_{0})=\mathcal{N}(\mathbf{x}_{t}\mid\sqrt{\alpha_{t}}\mathbf{x}_{0},(1-\alpha_{t})\mathbf{I})$, we can use the Tweedie’s formula (Efron,, 2011) to $p_{0\mid Y}(\mathbf{x}_{0}\mid\mathbf{y})$ and $p_{0}(\mathbf{x}_{0})$ to get
$\sqrt{\alpha_{t}}\hat{\mathbf{x}}_{0\mid t,Y}(\mathbf{x}_{t},\mathbf{y})$ $=\mathbf{x}_{t}+(1-\alpha_{t})\nabla_{\mathbf{x}_{t}}\log p_{t\mid Y}(\mathbf{x}_{t}\mid\mathbf{y})$ (9)
$\sqrt{\alpha_{t}}\hat{\mathbf{x}}_{0\mid t}(\mathbf{x}_{t})$ $=\mathbf{x}_{t}+(1-\alpha_{t})\nabla_{\mathbf{x}_{t}}\log p_{t}(\mathbf{x}_{t}).$ (10)
Since by Bayes theorem, $\nabla_{\mathbf{x}_{t}}\log p_{t\mid Y}(\mathbf{x}_{t}\mid\mathbf{y})=\nabla_{\mathbf{x}_{t}}\log p_{t}(\mathbf{x}_{t})+\log p_{Y\mid t}(\mathbf{y}\mid\mathbf{x}_{t})$, the results follows by simple algebra.
Since $p_{Y\mid 0}$ is Gaussian, using the DPS approximation on the second term in Eq. (8), we get
$\hat{\mathbf{x}}_{0\mid t,Y}(\mathbf{x}_{t},\mathbf{y})\approx\hat{\mathbf{x}}_{0\mid t}(\mathbf{x}_{t})-\left(\frac{1-\alpha_{t}}{\sqrt{\alpha_{t}}}\right)\frac{1}{2\sigma_{y}^{2}}\nabla_{\mathbf{x}_{t}}\lVert\mathbf{y}-\mathbf{A}(\hat{\mathbf{x}}_{0\mid t}(\mathbf{x}_{t}))\rVert_{2}^{2}.$ (11)
Replacing $1/2\sigma_{y}^{2}$ by $\eta$ to control the approximation error gives Eq. (6). In the following, we denote this approximation as
$\hat{\mathbf{x}}_{0\mid t,Y}(\mathbf{x}_{t},\mathbf{y})\approx\hat{\mathbf{x}}_{0\mid t}(\mathbf{x}_{t})-\left(\frac{1-\alpha_{t}}{\sqrt{\alpha_{t}}}\right)\eta\nabla_{\mathbf{x}_{t}}\lVert\mathbf{y}-\mathbf{A}(\hat{\mathbf{x}}_{0\mid t}(\mathbf{x}_{t}))\rVert_{2}^{2}\triangleq\tilde{\mathbf{x}}_{0\mid t}^{\eta}(\mathbf{x}_{t},\mathbf{y}).$ (12)
Proposition 3.
Assume that $r$ is a Lipschitz function that takes values in $[0,1]$. For any $\mathbf{x}_{t},\mathbf{y},\mathbf{s}$, the error in the value approximation $\hat{V}_{t}^{\tau}(\mathbf{x}_{t};\mathbf{s},\mathbf{y})=r(\tilde{\mathbf{x}}_{0\mid t}^{\eta}(\mathbf{x}_{t},\mathbf{y});\mathbf{s})/\tau$ with the true value $V_{t}^{\tau}(\mathbf{x}_{t};\mathbf{s},\mathbf{y})$ is bounded as
$|V_{t}^{\tau}(\mathbf{x}_{t};\mathbf{s},\mathbf{y})-\hat{V}_{t}^{\tau}(\mathbf{x}_{t};\mathbf{s},\mathbf{y})|\leq c_{\tau}c_{1}(t)+\frac{L_{r}}{\tau}\left(\sqrt{c_{2}(t)}+\sqrt{c_{3}(t)c_{4}^{\eta}(t)}\right),$ (13)
where $c_{1}(t)=\mathrm{Var}\left(r(X_{0};\mathbf{s})\mid\mathbf{x}_{t},\mathbf{y}\right)$, $c_{2}(t)=\mathrm{Var}(X_{0}\mid\mathbf{x}_{t},\mathbf{y})$, $c_{3}(t)=\mathrm{Var}(X_{0}\mid\mathbf{x}_{t})$,
$c_{4}^{\eta}(t)$ $=1+\mathrm{CV}^{2}(t)+\eta^{2}\lVert\mathbf{A}^{T}(\mathbf{y}-\mathbf{A}\hat{\mathbf{x}}_{0\mid t}(\mathbf{x}_{t}))\rVert_{\Sigma_{0\mid t}(\mathbf{x}_{t})}^{2}$ (14)
$\hskip 40.0pt-2\eta\langle\mathbf{A}(\hat{\mathbf{x}}_{0\mid t,Y}(\mathbf{x}_{t},\mathbf{y})-\hat{\mathbf{x}}_{0\mid t}(\mathbf{x}_{t})),\mathbf{y}-\mathbf{A}\hat{\mathbf{x}}_{0\mid t}(\mathbf{x}_{t})\rangle,$ (15)
where $\mathrm{CV}(t)\triangleq\frac{\sqrt{\mathrm{Var}(p_{Y\mid 0}(\mathbf{y}\mid X_{0})\mid\mathbf{x}_{t})}}{\mathbb{E}[p_{Y\mid 0}(\mathbf{y}\mid X_{0})\mid\mathbf{x}_{t}]}$ is the coefficient of variation of the likelihood function $p_{Y\mid 0}(\mathbf{y}\mid X_{0})$ given $\mathbf{x}_{t}$ and $c_{\tau}=e^{1/\tau}-1-1/\tau$ is a positive constant.
Remark 4.
In Proposition 3, the term $c_{1}(t)$ denotes the conditional variance of the reward given $\mathbf{x}_{t},\mathbf{y}$, $c_{2}(t)$ denotes the conditional variance of $X_{0}$ given $\mathbf{x}_{t},\mathbf{y}$, and $c_{3}(t)$ denotes the conditional variance of $X_{0}$ given only $\mathbf{x}_{t}$. Since the variance of the reverse distribution $p_{0\mid t}(\cdot\mid\mathbf{x}_{t})$ decreases as $t$ becomes smaller, we have that all the terms $c_{1}(t),c_{2}(t),c_{3}(t)$ are small when $t$ is small. Therefore, the approximation error is small when $t$ is small.
Proof. Since $r$ is a bounded random variable, assuming finite variance, we can use Bennett’s inequality for the log moment-generating function
$V_{t}^{\tau}(\mathbf{x}_{t};\mathbf{s},\mathbf{y})=\log\mathbb{E}[\exp(r(X_{0};\mathbf{s})/\tau)]\leq\frac{1}{\tau}\mathbb{E}[r(X_{0};\mathbf{s})]+c_{\tau}c_{1}(t).$ (16)
Then, we have
$|\log\mathbb{E}[\exp(r(X_{0};\mathbf{s})/\tau)]-\frac{1}{\tau}r(\tilde{\mathbf{x}}^{\eta}_{0\mid t,Y}(\mathbf{x}_{t},\mathbf{y}))|\leq\frac{1}{\tau}|\mathbb{E}[r(X_{0};\mathbf{s})]-r(\tilde{\mathbf{x}}^{\eta}_{0\mid t,Y}(\mathbf{x}_{t},\mathbf{y}))|+c_{\tau}c_{1}(t).$ (17)
Now, let us simplify the first term,
$|\mathbb{E}[r(X_{0};\mathbf{s})]$- $r(\tilde{\mathbf{x}}^{\eta}_{0\mid t,Y}(\mathbf{x}_{t},\mathbf{y});\mathbf{s})|\leq\mathbb{E}[|r(X_{0};\mathbf{s})-r(\tilde{\mathbf{x}}^{\eta}_{0\mid t,Y}(\mathbf{x}_{t},\mathbf{y});\mathbf{s})|]$ (18)
$\leq L_{r}\mathbb{E}[\lVert X_{0}-\tilde{\mathbf{x}}^{\eta}_{0\mid t,Y}(\mathbf{x}_{t},\mathbf{y})\rVert_{2}]$ (19)
$\leq L_{r}(\mathbb{E}[\lVert X_{0}-\hat{\mathbf{x}}_{0\mid t,Y}(\mathbf{x}_{t},\mathbf{y})\rVert_{2}]+\lVert\hat{\mathbf{x}}_{0\mid t,Y}(\mathbf{x}_{t},\mathbf{y})-\tilde{\mathbf{x}}^{\eta}_{0\mid t,Y}(\mathbf{x}_{t},\mathbf{y})\rVert_{2}).$ (20)
The first term can be bounded by $\sqrt{c_{2}(t)}$ using Cauchy-Schwarz inequality in $L^{2}$-probability space. For the second term, first we simplify
$\tilde{\mathbf{x}}^{\eta}_{0\mid t,Y}(\mathbf{x}_{t},\mathbf{y})=\hat{\mathbf{x}}_{0\mid t}(\mathbf{x}_{t})-\left(\frac{1-\alpha_{t}}{\sqrt{\alpha_{t}}}\right)\eta\nabla_{\mathbf{x}_{t}}\lVert\mathbf{y}-\mathbf{A}\hat{\mathbf{x}}_{0\mid t}(\mathbf{x}_{t})\rVert_{2}^{2},$
and then $\nabla_{\mathbf{x}_{t}}\lVert\mathbf{y}-\mathbf{A}\hat{\mathbf{x}}_{0\mid t}(\mathbf{x}_{t})\rVert_{2}^{2}=-(\nabla_{\mathbf{x}_{t}}\hat{\mathbf{x}}_{0\mid t}(\mathbf{x}_{t}))\mathbf{A}^{T}(\mathbf{y}-\mathbf{A}\hat{\mathbf{x}}_{0\mid t}(\mathbf{x}_{t}))$. Now,
$\nabla_{\mathbf{x}_{t}}\hat{\mathbf{x}}_{0\mid t}(\mathbf{x}_{t})$ $=\nabla_{\mathbf{x}_{t}}\int\mathbf{x}_{0}^{T}p_{0\mid t}(\mathbf{x}_{0}\mid\mathbf{x}_{t})\mathrm{d}\mathbf{x}_{0}=\int\nabla_{\mathbf{x}_{t}}p_{0\mid t}(\mathbf{x}_{0}\mid\mathbf{x}_{t})\mathbf{x}_{0}^{T}\mathrm{d}\mathbf{x}_{0}$ (21)
$=\int\nabla_{\mathbf{x}_{t}}\log p_{0\mid t}(\mathbf{x}_{0}\mid\mathbf{x}_{t})\mathbf{x}_{0}^{T}p_{0\mid t}(\mathbf{x}_{0}\mid\mathbf{x}_{t})\mathrm{d}\mathbf{x}_{0}.$ (22)
Now, we shall compute
$\nabla_{\mathbf{x}_{t}}\log p_{0\mid t}(\mathbf{x}_{0}\mid\mathbf{x}_{t})=\nabla_{\mathbf{x}_{t}}\log p_{t\mid 0}(\mathbf{x}_{t}\mid\mathbf{x}_{0})-\nabla_{\mathbf{x}_{t}}\log p_{t}(\mathbf{x}_{t}).$ (23)
But since $\sqrt{\alpha_{t}}\hat{\mathbf{x}}_{0\mid t}(\mathbf{x}_{t})=\mathbf{x}_{t}+(1-\alpha_{t})\nabla_{\mathbf{x}_{t}}\log p_{t}(\mathbf{x}_{t})$, and $\nabla_{\mathbf{x}_{t}}\log p_{t\mid 0}(\mathbf{x}_{t}\mid\mathbf{x}_{0})=\frac{1}{1-\alpha_{t}}(\sqrt{\alpha_{t}}\mathbf{x}_{0}-\mathbf{x}_{t})$, which gives
$\left(\frac{1-\alpha_{t}}{\sqrt{\alpha_{t}}}\right)\nabla_{\mathbf{x}_{t}}\log p_{0\mid t}(\mathbf{x}_{0}\mid\mathbf{x}_{t})=\mathbf{x}_{0}-\hat{\mathbf{x}}_{0\mid t}(\mathbf{x}_{t}),$ (24)
which gives
$\left(\frac{1-\alpha_{t}}{\sqrt{\alpha_{t}}}\right)\nabla_{\mathbf{x}_{t}}\hat{\mathbf{x}}_{0\mid t}(\mathbf{x}_{t})$ $=\mathbb{E}_{X_{0}\sim p_{0\mid t}(\mathbf{x}_{t})}[(X_{0}-\hat{\mathbf{x}}_{0\mid t}(\mathbf{x}_{t}))X_{0}^{T}]$ (25)
$=\mathbb{E}_{p_{0\mid t}(\mathbf{x}_{t})}[X_{0}X_{0}^{T}]-\mathbb{E}_{p_{0\mid t}(\mathbf{x}_{t})}[X_{0}]\mathbb{E}_{p_{0\mid t}(\mathbf{x}_{t})}[X_{0}^{T}]$ (26)
$=\mathbb{E}_{p_{0\mid t}(\mathbf{x}_{t})}[(X_{0}-\hat{\mathbf{x}}_{0\mid t}(\mathbf{x}_{t}))(X_{0}-\hat{\mathbf{x}}_{0\mid t}(\mathbf{x}_{t}))^{T}]\triangleq\Sigma_{0\mid t}(\mathbf{x}_{t}),$ (27)
which is precisely the covariance matrix of $X_{0}$ given $\mathbf{x}_{t}$.
Thus, we finally get,
$\tilde{\mathbf{x}}^{\eta}_{0\mid t,Y}(\mathbf{x}_{t},\mathbf{y})$ $=\hat{\mathbf{x}}_{0\mid t}(\mathbf{x}_{t})+\eta\Sigma_{0\mid t}(\mathbf{x}_{t})\mathbf{A}^{T}(\mathbf{y}-\mathbf{A}\hat{\mathbf{x}}_{0\mid t}(\mathbf{x}_{t}))$ (28)
Next, note that since $\mathbb{E}_{p_{0\mid t}(\mathbf{x}_{t})}[p_{Y\mid 0}(\mathbf{y}\mid X_{0})]=p_{Y\mid t}(\mathbf{y}\mid\mathbf{x}_{t})$, we can define $f(X_{0})=\frac{p_{Y\mid 0}(\mathbf{y}\mid X_{0})}{\mathbb{E}_{p_{0\mid t}(\mathbf{x}_{t})}[p_{Y\mid 0}(\mathbf{y}\mid X_{0})]}$, whose expectation is $\mathbb{E}_{p_{0\mid t}(\mathbf{x}_{t})}[f(X_{0})]=1$. Further, it is easy to see that $\hat{\mathbf{x}}_{0\mid t,Y}(\mathbf{x}_{t},\mathbf{y})=\mathbb{E}_{p_{0\mid t}(\mathbf{x}_{t})}[X_{0}f(X_{0})]$. Now, we are ready to bound the final term as follows
$\lVert\hat{\mathbf{x}}_{0\mid t,Y}$ $(\mathbf{x}_{t},\mathbf{y})-\tilde{\mathbf{x}}^{\eta}_{0\mid t,Y}(\mathbf{x}_{t},\mathbf{y})\rVert_{2}$ (29)
$=\lVert\mathbb{E}_{p_{0\mid t}(\mathbf{x}_{t})}[X_{0}f(X_{0})]-\hat{\mathbf{x}}_{0\mid t}(\mathbf{x}_{t})-\eta\Sigma_{0\mid t}(\mathbf{x}_{t})\mathbf{A}^{T}(\mathbf{y}-\mathbf{A}\hat{\mathbf{x}}_{0\mid t}(\mathbf{x}_{t}))\rVert_{2}$ (30)
$=\lVert\mathbb{E}_{p_{0\mid t}(\mathbf{x}_{t})}[X_{0}f(X_{0})-\hat{\mathbf{x}}_{0\mid t}(\mathbf{x}_{t})f(X_{0})]$ (31)
$\hskip 40.0pt-\eta\mathbb{E}_{p_{0\mid t}(\mathbf{x}_{t})}[(X_{0}-\hat{\mathbf{x}}_{0\mid t}(\mathbf{x}_{t}))(X_{0}-\hat{\mathbf{x}}_{0\mid t}(\mathbf{x}_{t}))^{T}]\mathbf{A}^{T}(\mathbf{y}-\mathbf{A}\hat{\mathbf{x}}_{0\mid t}(\mathbf{x}_{t}))\rVert_{2}$ (32)
$=\lVert\mathbb{E}_{p_{0\mid t}(\mathbf{x}_{t})}[(X_{0}-\hat{\mathbf{x}}_{0\mid t}(\mathbf{x}_{t}))(f(X_{0})-\eta(X_{0}-\hat{\mathbf{x}}_{0\mid t}(\mathbf{x}_{t}))^{T}\mathbf{A}^{T}(\mathbf{y}-\mathbf{A}\hat{\mathbf{x}}_{0\mid t}(\mathbf{x}_{t}))))]\rVert_{2}$ (33)
$\leq\mathbb{E}_{p_{0\mid t}(\mathbf{x}_{t})}[\lVert(X_{0}-\hat{\mathbf{x}}_{0\mid t}(\mathbf{x}_{t}))(f(X_{0})-\eta(X_{0}-\hat{\mathbf{x}}_{0\mid t}(\mathbf{x}_{t}))^{T}\mathbf{A}^{T}(\mathbf{y}-\mathbf{A}\hat{\mathbf{x}}_{0\mid t}(\mathbf{x}_{t}))))\rVert_{2}]$ (34)
$=\mathbb{E}_{p_{0\mid t}(\mathbf{x}_{t})}[\lVert X_{0}-\hat{\mathbf{x}}_{0\mid t}(\mathbf{x}_{t})\rVert_{2}|f(X_{0})-\eta(X_{0}-\hat{\mathbf{x}}_{0\mid t}(\mathbf{x}_{t}))^{T}\mathbf{A}^{T}(\mathbf{y}-\mathbf{A}\hat{\mathbf{x}}_{0\mid t}(\mathbf{x}_{t}))|]$ (35)
$\leq\sqrt{c_{3}(t)}\sqrt{c_{4}^{\eta}(t)},$ (36)
where the last step follows by Cauchy-Schwarz inequality in $L^{2}$ probability space, where
$c_{3}(t)\triangleq\mathbb{E}_{p_{0\mid t}(\mathbf{x}_{t})}[\lVert X_{0}-\mathbb{E}_{p_{0\mid t}(\mathbf{x}_{t})[X_{0}]}\rVert_{2}^{2}]=\mathrm{Var}(X_{0}\mid\mathbf{x}_{t}).$ (37)
and at last, we have
$c_{4}^{\eta}(t)$ $\triangleq\mathbb{E}_{p_{0\mid t}(\mathbf{x}_{t})}[(f(X_{0})-\eta(X_{0}-\hat{\mathbf{x}}_{0\mid t}(\mathbf{x}_{t}))^{T}\mathbf{A}^{T}(\mathbf{y}-\mathbf{A}\hat{\mathbf{x}}_{0\mid t}(\mathbf{x}_{t})))^{2}]$ (38)
$=\mathbb{E}_{p_{0\mid t}(\mathbf{x}_{t})}[f(X_{0})^{2}]+\eta^{2}\mathbb{E}[((X_{0}-\hat{\mathbf{x}}_{0\mid t}(\mathbf{x}_{t}))^{T}\mathbf{A}^{T}(\mathbf{y}-\mathbf{A}\hat{\mathbf{x}}_{0\mid t}(\mathbf{x}_{t})))^{2}]$ (39)
$\hskip 40.0pt-2\eta\mathbb{E}_{p_{0\mid t}(\mathbf{x}_{t})}[(X_{0}f(X_{0})-f(X_{0})\hat{\mathbf{x}}_{0\mid t}(t))^{T}]\mathbf{A}^{T}(\mathbf{y}-\mathbf{A}\hat{\mathbf{x}}_{0\mid t}(\mathbf{x}_{t}))$ (40)
$=\mathbb{E}_{p_{0\mid t}(\mathbf{x}_{t})}[f(X_{0})^{2}]+\eta^{2}\lVert\mathbf{A}^{T}(\mathbf{y}-\mathbf{A}\hat{\mathbf{x}}_{0\mid t}(\mathbf{x}_{t}))\rVert_{\Sigma_{0\mid t}(\mathbf{x}_{t})}^{2}$ (41)
$\hskip 40.0pt-2\eta\langle\mathbf{A}(\hat{\mathbf{x}}_{0\mid t,Y}(\mathbf{x}_{t},\mathbf{y})-\hat{\mathbf{x}}_{0\mid t}(\mathbf{x}_{t})),\mathbf{y}-\mathbf{A}\hat{\mathbf{x}}_{0\mid t}(\mathbf{x}_{t})\rangle.$ (42)
Since $\mathbb{E}_{p_{0\mid t}(\mathbf{x}_{t})}[f(X_{0})^{2}]=\frac{\mathbb{E}_{p_{0\mid t}(\mathbf{x}_{t})}[(p_{Y\mid 0}(\mathbf{y}\mid X_{0}))^{2}]}{(\mathbb{E}_{p_{0\mid t}(\mathbf{x}_{t})}[p_{Y\mid 0}(\mathbf{y}\mid X_{0})])^{2}}=1+\mathrm{CV}^{2}(t)$, where $\mathrm{CV}(t)=\frac{\sqrt{\mathrm{Var}(p_{Y\mid 0}(\mathbf{y}\mid X_{0})\mid\mathbf{x}_{t})}}{\mathbb{E}[p_{Y\mid 0}(\mathbf{y}\mid X_{0})\mid\mathbf{x}_{t}]}$ is the coefficient of variation of the likelihood function given $\mathbf{x}_{t}$.
A.3 Algorithm
We summarize our framework in Algorithm 1. The GroupSize step in line $5$ computes the group-size at time $t$ and can be changed to obtain various search strategies: Best-of-N, Greedy Search, and RFJ Search, and the Resample step in line $6$ samples the indices within the groups as described in the main paper. Line $3$ is specific to the diffusion samplers and how they implement it.
Here, we roughly explain how the entire algorithm is implemented in all the three diffusion samplers.
DPS: Compute the denoised mean and the clean data estimate $\boldsymbol{\mu}_{t}[i],\hat{\mathbf{x}}_{0\mid t}[i]$ from $\mathbf{x}_{t+1}[i]$. Compute $\mathbf{g}_{t}[i]=\nabla_{\mathbf{x}_{t}}\lVert\mathbf{y}-\mathbf{A}\hat{\mathbf{x}}_{0\mid t}(\mathbf{x}_{t+1}[i])\rVert_{2}^{2}$ and use it to update $\hat{\mathbf{x}}_{0\mid t,Y}[i]\leftarrow\hat{\mathbf{x}}_{0\mid t}[i]-\frac{1-\alpha_{t}}{\sqrt{\alpha_{t}}}\eta\mathbf{g}_{t}[i]$, and $\boldsymbol{\mu}_{t}[i]\leftarrow\boldsymbol{\mu}_{t}[i]-\zeta\mathbf{g}_{t}[i]$. Then, compute rewards based on $\hat{\mathbf{x}}_{0\mid t,Y}[i]$ to resample promising indices. Take the reverse diffusion step on the resampled $\boldsymbol{\mu}_{t}[I[i]]$ to obtain $\mathbf{x}_{t}[i]$.
DAPS: Compute the clean data estimate $\hat{\mathbf{x}}_{0\mid t}[i]$ from $\mathbf{x}_{t+1}[i]$. Compute the rewards based on $\hat{\mathbf{x}}_{0\mid t}[i]$, resample, and then take MCMC steps, starting from the resampled particles, to perform a local Langevin sampling (Zhang et al.,, 2024). In the end, we obtain $\hat{\mathbf{x}}_{t,Y}[i]$, from which we sample $\mathbf{x}_{t}[i]$ by adding appropriate decoupled noise (Zhang et al.,, 2024).
MPGD: Compute the clean data estimate $\hat{\mathbf{x}}_{0\mid t}[i]$ from $\mathbf{x}_{t+1}[i]$. Compute gradient $\mathbf{g}_{t}=\nabla_{\mathbf{x}_{0\mid t}}\lVert\mathbf{y}-\mathbf{A}\hat{\mathbf{x}}_{0\mid t}[i]\rVert_{2}^{2}$. Take $\hat{\mathbf{x}}_{0\mid t,Y}[i]\leftarrow\hat{\mathbf{x}}_{0\mid t}[i]-\frac{1-\alpha_{t}}{\sqrt{\alpha_{t}}}\eta\mathbf{g}_{t}[i]$, and $\hat{\mathbf{x}}_{0\mid t}[i]\leftarrow\hat{\mathbf{x}}_{0\mid t}[i]-\zeta\mathbf{g}_{t}$. Resample indices based on rewards computed from $\hat{\mathbf{x}}_{0\mid t,Y}[i]$. Then, using the particles corresponding to the sampled indices $\hat{\mathbf{x}}_{0\mid t,Y}[I[i]]$, take a reverse DDIM step (He et al.,, 2024; Song et al., 2021a, ).
Algorithm 1 Inference-Time Search with Side Information for Inverse Problems
1:Side information $\mathbf{s}$, observation $\mathbf{y}$, reward function $r$, resampling parameter $B$, number of particles NN, temperature $\tau>0$
2:Initialize NN particles: $\mathbf{x}_{T}[i]\sim\mathcal{N}(0,\mathbf{I})$ for $1\leq i\leq N$
3:for t=T1t=T-1 to 0 do
4:  Sample $\mathbf{x}_{t}[i]\sim p_{t\mid t+1,Y}(\cdot\mid\mathbf{x}_{t+1}[i],\mathbf{y}),$ $\triangleright$ Sample candidate particles
5:  $r[i]\leftarrow r(\hat{\mathbf{x}}_{0\mid t,Y}[i];\mathbf{s})$ $\triangleright$ Compute reward using side information
6:  $g_{t}\leftarrow\textsc{GroupSize}(N,B,t)$ $\triangleright$ Compute the group size at step $t$ for resampling
7:  $I\leftarrow\textsc{Resample}(r,g_{t},\tau)$ $\triangleright$ Resample indices with replacement among the groups of size $g_{t}$
8:  $\mathbf{x}_{t}[i]\leftarrow\mathbf{x}_{t}[I[i]]$ $\triangleright$ Retain the particles in the resampled indices
9:Select $\mathbf{x}^{*}_{0}$ from $\hat{\mathbf{x}}_{0\mid 0}[1:N]$ (e.g., via reward maximization) return $\mathbf{x}^{*}_{0}$
Appendix B Additional Experiments
B.1 DPS
We present additional qualitative examples for our DPS experiments in Figures 8 and 9, corresponding to settings that use face and text as side information, respectively.
Side Measurement DPS RGG RFJS (Ours) Ground Truth Side Measurement DPS RGG RFJS (Ours) Ground Truth Box Inpainting Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Box Inpainting Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Super Resolution Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Super Resolution Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Motion Deblur Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Motion Deblur Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Gaussian Deblur Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Gaussian Deblur Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Nonlinear Deblur Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Nonlinear Deblur Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Blind Deblur Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Blind Deblur Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Figure 8: Additional samples using DPS as the base sampler.
Measurement DPS RFJ Search Ground Truth Measurement DPS RFJ Search Ground Truth Box Inpainting (138) Refer to caption Refer to caption Refer to caption Refer to caption Super-resolution (32) Refer to caption Refer to caption Refer to caption Refer to caption Side: An elephant stands in tall grass on a savanna with distant trees. Side: A fluffy gray kitten with white markings. Motion Deblur Refer to caption Refer to caption Refer to caption Refer to caption Gaussian Deblur Refer to caption Refer to caption Refer to caption Refer to caption Side: A reddish-brown golden retriever stands upright in a forest, green leaves on the left. Side: A wet golden retriever stands on a sandy beach, facing left. Nonlinear Deblur Refer to caption Refer to caption Refer to caption Refer to caption Blind Deblur Refer to caption Refer to caption Refer to caption Refer to caption Side: A male lion with a black-to-rust mane, surrounded by soft dry grass. Side: A tabby cat with dark stripes.
Figure 9: Qualitative comparison on ImageNet with textual side information. For highly degraded observations, DPS and BlindDPS often produce artifacts, whereas our method reduces these by better aligning reconstructions with the description.
B.2 DAPS
Setup: We employ the search and gradient modules to infuse side information using DAPS as the base sampler. We consider two challenging tasks, box inpainting with a box of size $96\times 96$ and super-resolution with downsampling factor of $10$. For gradient guidance, we used a scale of $15$ with respect to the noise being added to $\hat{\mathbf{x}}_{0\mid t}(\mathbf{x}_{t})$ after MCMC steps (Zhang et al.,, 2024). DAPS uses fewer diffusion steps $(200)$ than DPS $(1000)$. Further, the algorithm is based on a graphical model that allows for more inherent exploration due to the decoupling between consecutive steps. Therefore, for search algorithms, relatively smaller bases are preferable, and hence $B=4$ is chosen.
Results: The qualitative results are given in Figure 10, while the quantitative metrics are given in 3. Observe that for the task of inpainting, our search algorithm shows a significant improvement over DAPS, particularly in the FaceSimilarity metric.
 Box Inpainting  Super Resolution ($\times 10$)
Method FaceSimilarity ($\downarrow$) PSNR ($\uparrow$) LPIPS ($\downarrow$) SSIM ($\uparrow$) FaceSimilarity ($\downarrow$) PSNR ($\uparrow$) LPIPS ($\downarrow$) SSIM ($\uparrow$)
RFJS (ours) $\textbf{0.423}\scriptstyle\pm 0.10$ $\textbf{28.720}\scriptstyle\pm 1.35$ $\textbf{0.140}\scriptstyle\pm 0.03$ $\textbf{0.788}\scriptstyle\pm 0.03$ $\underline{0.654}\scriptstyle\pm 0.11$ $\underline{25.228}\scriptstyle\pm 1.34$ $\textbf{0.282}\scriptstyle\pm 0.03$ $\underline{0.661}\scriptstyle\pm 0.04$
GS (ours) $0.511\scriptstyle\pm 0.12$ $28.640\scriptstyle\pm 1.43$ $\underline{0.140}\scriptstyle\pm 0.03$ $\underline{0.787}\scriptstyle\pm 0.03$ $0.760\scriptstyle\pm 0.12$ $\textbf{25.271}\scriptstyle\pm 1.36$ $0.285\scriptstyle\pm 0.03$ $\textbf{0.662}\scriptstyle\pm 0.04$
RGG $\underline{0.436}\scriptstyle\pm 0.12$ $28.410\scriptstyle\pm 1.39$ $0.141\scriptstyle\pm 0.03$ $0.784\scriptstyle\pm 0.03$ $\textbf{0.579}\scriptstyle\pm 0.13$ $25.210\scriptstyle\pm 1.34$ $\underline{0.282}\scriptstyle\pm 0.03$ $0.659\scriptstyle\pm 0.04$
BON $0.611\scriptstyle\pm 0.14$ $\underline{28.660}\scriptstyle\pm 1.45$ $0.141\scriptstyle\pm 0.03$ $0.787\scriptstyle\pm 0.03$ $0.909\scriptstyle\pm 0.11$ $25.220\scriptstyle\pm 1.38$ $0.285\scriptstyle\pm 0.03$ $0.660\scriptstyle\pm 0.04$
DAPS $0.739\scriptstyle\pm 0.18$ $28.290\scriptstyle\pm 1.53$ $0.142\scriptstyle\pm 0.03$ $0.784\scriptstyle\pm 0.03$ $1.020\scriptstyle\pm 0.14$ $25.170\scriptstyle\pm 1.35$ $0.285\scriptstyle\pm 0.03$ $0.659\scriptstyle\pm 0.04$
Table 3: Comparison of metrics for various inverse problems using DAPS as the base sampler. For each metric, the best result is shown in bold, and the second best is underlined. We observe that our RFJ Search-based algorithm has the best or the second-best performance in all the tasks.
Side Measurement DAPS RGG RFJS (Ours) Ground Truth Side Measurement DAPS Gradient RFJS (Ours) Ground Truth Box Inpainting (96) Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Super-resolution (10) Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Figure 10: Qualitative comparison of algorithms using DAPS as the base sampler. Our method offers better reconstructions aligned with the identity.
B.3 MPGD
Setup: As in the paper He et al., (2024), we choose super-resolution and Gaussian deblurring as the tasks, along with additional task of box inpainting. For box inpainting, we used a box of size $64\times 64$ at the center of the face. Further, we modified the down-sampling scale of super-resolution from $4$ to $6$, and the intensity of the kernel in Gaussian deblur from $3$ to $5$ to make the tasks more challenging. Even though MPGD uses $100$ DDIM steps in generation, its exploratory capabilities are similar to DPS. Therefore, we cannot use very large base $B$, whence, we choose $B=8$ for box inpainting, and super-resolution. For Gaussian deblur, we found that using $B=8$ becomes detrimental for other metrics, and so $B=16$ is used. For the gradient guidance, the scales of $0.5$ for box inpainting, and $0.25$ for super-resolution and Gaussian deblur are chosen carefully to avoid overfitting. This is the scale relative to the gradient with respect to the measurement. For more discussion and examples on the effect of gradient scale, see Appendix B.5.
Results: Experimental results with MPGD (He et al.,, 2024) as baseline algorithm are given in Table 4. We observe that using RFJ Search significantly enhances the FaceSimilarity (FS) metric, while improving the other metrics. The reconstructions that utilize the side information exhibit strong identity match with the one described by the measurement, which is reflected in the FS metric.
Sampler Task Method FaceSimilarity ($\downarrow$) PSNR ($\uparrow$) LPIPS ($\downarrow$) SSIM ($\uparrow$)
MPGD  Box Inpainting  (64) RFJS (8) (ours) $\mathbf{0.542}\scriptstyle\pm 0.08$ $\mathbf{29.81}\scriptstyle\pm 1.44$ $\mathbf{0.102}\scriptstyle\pm 0.02$ $\mathbf{0.852}\scriptstyle\pm 0.02$
GS (8) (ours) $\underline{0.587}\scriptstyle\pm 0.10$ $\underline{29.44}\scriptstyle\pm 1.75$ $\underline{0.102}\scriptstyle\pm 0.02$ $\underline{0.851}\scriptstyle\pm 0.02$
RGG (0.5) $0.609\scriptstyle\pm 0.08$ $29.24\scriptstyle\pm 1.30$ $0.103\scriptstyle\pm 0.02$ $0.850\scriptstyle\pm 0.02$
BON $0.661\scriptstyle\pm 0.08$ $29.35\scriptstyle\pm 1.82$ $0.102\scriptstyle\pm 0.02$ $0.851\scriptstyle\pm 0.02$
MPGD $0.766\scriptstyle\pm 0.07$ $29.09\scriptstyle\pm 1.27$ $0.103\scriptstyle\pm 0.02$ $0.848\scriptstyle\pm 0.02$
 Super Resolution (6) RFJS (8) (ours) $\mathbf{0.834}\scriptstyle\pm 0.09$ $\mathbf{24.50}\scriptstyle\pm 1.48$ $\mathbf{0.242}\scriptstyle\pm 0.04$ $\mathbf{0.666}\scriptstyle\pm 0.06$
GS (8) (ours) $0.878\scriptstyle\pm 0.08$ $\underline{24.45}\scriptstyle\pm 1.47$ $0.247\scriptstyle\pm 0.03$ $0.660\scriptstyle\pm 0.06$
RGG (0.25) $\underline{0.854}\scriptstyle\pm 0.07$ $24.39\scriptstyle\pm 1.44$ $0.246\scriptstyle\pm 0.03$ $0.656\scriptstyle\pm 0.05$
BON $0.964\scriptstyle\pm 0.09$ $24.44\scriptstyle\pm 1.58$ $\underline{0.244}\scriptstyle\pm 0.04$ $\underline{0.664}\scriptstyle\pm 0.06$
MPGD $1.037\scriptstyle\pm 0.07$ $24.39\scriptstyle\pm 1.45$ $0.249\scriptstyle\pm 0.03$ $0.657\scriptstyle\pm 0.06$
 Gaussian Deblur (5) RFJS (16) (ours) $\underline{0.848}\scriptstyle\pm 0.07$ $\underline{24.19}\scriptstyle\pm 1.40$ $\mathbf{0.229}\scriptstyle\pm 0.03$ $\underline{0.638}\scriptstyle\pm 0.06$
GS (16) (ours) $0.893\scriptstyle\pm 0.07$ $24.14\scriptstyle\pm 1.39$ $\underline{0.233}\scriptstyle\pm 0.03$ $0.637\scriptstyle\pm 0.06$
RGG (0.25) $\mathbf{0.846}\scriptstyle\pm 0.05$ $24.11\scriptstyle\pm 1.34$ $0.235\scriptstyle\pm 0.03$ $0.634\scriptstyle\pm 0.05$
BON $0.950\scriptstyle\pm 0.07$ $\mathbf{24.20}\scriptstyle\pm 1.38$ $0.233\scriptstyle\pm 0.03$ $\mathbf{0.640}\scriptstyle\pm 0.06$
MPGD $1.026\scriptstyle\pm 0.06$ $24.09\scriptstyle\pm 1.35$ $0.236\scriptstyle\pm 0.03$ $0.634\scriptstyle\pm 0.06$
Table 4: Quantitative comparison of reconstruction metrics in case of inverse problems with MPGD as the base sampler. For each metric, the best result is shown in bold, and the second best is underlined. Observe that our RFJSsearch algorithm has the best or the second-best performance in all the tasks. In Gaussian deblur, our search algorithm is only marginally worse than the best metrics attained. The value in the brackets indicates the resampling rate for search algorithms, and gradient scale for Gradient algorithm.
Side Measurement MPGD RGG RFJS (Ours) Ground Truth Side Measurement MPGD RGG RFJS (Ours) Ground Truth Box Inpainting Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Box Inpainting Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Super Resolution Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Super Resolution Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Gaussian Deblur Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Gaussian Deblur Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Figure 11: Qualitative comparison of algorithms using MPGD as the base sampler. Our method offers better reconstructions aligned with the identity. Notice that while the identity is preserved, the exact reconstruction might not be possible as witnessed in, for example, Box inpainting. The ground truth has a smiling face whereas the reconstruction does not, although being the same person. Thus, PSNR improvements over the base sampler might be small, but FaceSimilarity improvements are significant.
B.4 Evaluating Search Strategies in a 2D Setting
To illustrate the effect of side information in inference-time search, we consider a simple 2D setup. The prior on $x_{0}$ is a mixture of Gaussians, shown in the leftmost image of Figure 13. A ground-truth sample $x_{0}$ is drawn from this prior, and a random forward operator $A$ is generated with fixed norm and random orientation. The observation is then $y=Ax_{0}+n$. The corresponding posterior $p(x_{0}\mid y)$, approximated by DPS, is also shown in Figure 13. Because this is an ill-posed inverse problem, the posterior is multimodal and DPS fails to reconstruct the correct solution.
We then add side information of the form $s=Bx_{0}+n_{s}$, where $B$ is chosen so that $Bx_{0}$ is orthogonal to $Ax_{0}$. The reward function is defined as $r(x_{t},s)=-\lVert s-B\hat{x}_{0}(x_{t})\rVert_{2}^{2}$, where $\hat{x}_{0}(x_{t})$ is the Tweedie estimate of the clean signal from the noisy state $x_{t}$. Using this reward, RFJ Search produces posterior samples $p(x_{0}\mid y,s)$ as shown in the three rightmost images of Figure 13. As the resampling base $B$ decreases, the particles concentrate more tightly around the ground truth, highlighting how side information guides inference.
Figure 12: Comparison of performance of RFJ Search (RFJS) and Greedy Search (GS) as a function of $B$ for a randomly generated reward network $r_{\theta}$. RFJS outperforms GS across all values of $B$.
To compare RFJ and Greedy Search, we next consider a more realistic scenario where side information is generated through a neural network reward model. For example, in face reconstruction tasks, side information may come from the embedding of an additional image, while for text-conditioned tasks, it may come from a text encoder. To mimic this, in each trial we sample a ground truth $x_{0}$, a forward operator $A$, and a reward network $r_{\theta}$ with randomly initialized weights. The side information is defined as $s=r_{\theta}(x_{0})+n_{s}$, and the reward is the cosine similarity between $r_{\theta}(\hat{x}_{0}(x_{t}))$ and $s$. We repeat this experiment 16,000 times, and for each trial run both RFJ and Greedy Search 128 times with $N=8$ particles, varying the resampling base $B$ over powers of two from $1$ to $256$. The average PSNR values are reported in Figure 12. Across all settings, RFJ Search consistently outperforms Greedy Search, with the best performance observed at $B=4$.
Figure 13: Illustration of the effect of $B$ in utilization of side information for the reconstruction in a linear inverse problem with a mixture of Gaussian prior.
Figure 14: Effect of reward-gradient guidance in diffusion-based inverse problems. We show 4 runs with different random seeds (rows), and for each seed we vary the gradient scale across 4 settings (columns). Within each row, the noise realization is identical and only the gradient scale changes; within each column, the gradient scale is fixed while the random seed varies. The ground truth and degraded input are the same for all reconstructions. This arrangement reveals two key observations: (1) The reward gradient influences fine details, such as wrinkles and facial lines without altering the overall facial structure; the structure is primarily determined by the initial noise realization. (2) Different seeds reconstruct different face structures, highlighting the multi-modal nature of the problem. This demonstrates why using multiple particles and performing search across them is beneficial: it enables exploration of structurally different hypotheses while the reward gradient refines locally.
DPS Gradient (1.6) Ground Truth Super Resolution (4) Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Figure 15: Qualitative comparison of the effect of gradient scale on reconstruction paths.
B.5 Experiments on Effect of Gradient Scale
Gradient Guidance Limitations. While guiding the reverse diffusion process with reward gradients can help generate images with higher reward scores, this approach has several limitations. First, as shown in Figures 14, 15, gradient-based guidance primarily adds fine details, such as wrinkles or texture, to the reconstruction, but it cannot significantly alter the global structure of the face. To isolate the effect of the gradient, we used fixed noise realizations for both the gradient-based and baseline methods. The results show that changes are mostly confined to local details, implying that if the sampling trajectory is poor, gradient guidance alone cannot compensate. This highlights the need for search-based methods that can explore a wider range of trajectories during inference.
Second, this method is sensitive to the choice of gradient scale. In the visual examples, we used a relatively large scale of 1.6 to make the gradient’s effect more visible; however, such high scales often degrade other metrics like PSNR and SSIM and introduce artifacts. Empirically, we found that a scale around $0.5$ yields the best balance when the base sampler is DPS or BlindDPS, consistently improving the FaceSimilarity metric while preserving other evaluation metrics and avoiding artifacts (see Figure 15). Moreover, the sensitivity to gradient scale increases when the number of reverse diffusion steps is small. For instance, in DAPS (Figure 14) and MPGD, where the number of steps is limited to 200 and 100 respectively, larger scales quickly lead to visible artifacts, as demonstrated in Figures 16, and 17.
It is well established in deep learning research that deep (convolutional) neural networks are vulnerable to gradient-based adversarial attacks (Goodfellow et al.,, 2015). Consequently, using such networks to provide reward-based guidance inherits these vulnerabilities. However, when combined with diffusion samplers, this susceptibility is partially alleviated, as the diffusion process can help steer trajectories away from adversarially induced local minima. This mitigating effect is particularly evident when using a large number of sampling steps (e.g., 1000 steps in DPS). In contrast, samplers with fewer steps (e.g., 100 steps in MPGD) exhibit increased sensitivity to the gradient scale, as illustrated in Figures 16 and 17.
MPGD Gradient (0.5) Gradient (1.0) Gradient (1.5) Ground Truth MPGD Gradient (0.5) Gradient (1.0) Gradient (1.5) Ground Truth Box Inpainting (64) Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Figure 16: Qualitative comparison of the effect of gradient scale on reconstruction paths. Notice that while the base reconstruction is reasonable, adding the gradient can degrade it if the scale is very large. The final scale used in the experiments is $0.5$.
MPGD Gradient (0.25) Gradient (0.5) Gradient (1.0) Ground Truth MPGD Gradient (0.25) Gradient (0.5) Gradient (1.0) Ground Truth Super Resolution (6) Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Figure 17: Qualitative comparison of the effect of gradient scale on reconstruction paths for super resolution task. Notice that while the base reconstruction is reasonable, adding the gradient can degrade it if the scale is very large. The final scale used in the experiments is $0.25$.
B.6 Effect of Number of Particles
Figure 18: Scaling of search algorithms with respect to the number of particles.
Although we do not explicitly define a second base, the recursive structure of our RFJS algorithm implicitly induces a small base, denoted by $b$, which refers to the interval at which the smallest groups are resampled. For instance, with 8 particles and a base $B=16$, the smallest groups of size 2 are resampled every 4 steps, so the induced small base is $b=4$. This hierarchical resampling schedule enables finer control over exploration without over-committing the entire particle set too aggressively. For analyzing the effect of particle count in RFJ Search (see Figure 19), we fix the small base $b$ so that groups of the same size follow the same resampling dynamics, even across experiments with different total numbers of particles. This ensures that any differences we observe are attributable to the number of particles, not to changes in resampling frequency. On the other hand, when comparing RFJ Search vs Greedy Search, we fix the big base $B$ to ensure both methods resample at the same overall frequency. This is why, in our plots, the $B$ values for Greedy search vary with the number of particles (e.g., $B=4,8,16$ for $2,4,8$ particles respectively), matching the corresponding configurations in RFJ Search.
(a) Nonlinear Deblur
(b) Super Resolution (4)
(c) Gaussian Deblur
(d) Motion Deblur
Figure 19: Effect of the number of particles NN on the FaceSimilarity metric in DPS. As NN increases, we proportionally increase the base to encourage greater exploration. Specifically, for $N=2,4,8$, the corresponding base values are $B=4,8,16$. Using many particles with a small value for base $B$ can overly emphasize reward maximization, which may degrade other evaluation metrics.
(a) Box Inpainting (Box: 64)
(b) Super-resolution (Scale: 6)
(c) Gaussian Blur (Intensity: 5)
Figure 20: Effect of number of particles NN on the FaceSimilarity metric. RFJ Search algorithm offers the best scaling performance, followed by Greedy Search algorithm. Finally, BestOfN performance improves, but only marginally. $64$ indicates the size of the box for inpainting, 66 indicates the down-sampling factor in super-resolution, and $5$ is the intensity of the Gaussian kernel in Gaussian deblur. The base resampling rate $B$ is $8$ for both the search algorithms in box inpainting and super-resolution whereas $B$ is $16$ for the Gaussian deblur.
B.7 Runtimes
[…]
Particles BON Greedy Search RFJ Search
1 55 - -
2 65 75 75
4 102 118 131
8 180 195 241
(a) Task: Box inpainting. Baseline: DPS
Particles BON Greedy Search RFJS
1 61 - -
2 72 91 91
4 125 141 157
8 229 245 290
(b) Task: Box inpainting. Baseline: DAPS
Particles BON Greedy Search RFJS
1 3 - -
2 4 5 5
4 5 8 9
8 8 12 23
(c) Task: Box inpainting. Baseline: MPGD
Table 5: Runtime (seconds) vs number of particles for BON, Greedy Search, and RFJ Search methods on DPS, DAPS and MPGD $(B=8)$. The baseline algorithm corresponds to $N=1$.
Appendix C Hyperparameters
In our implementation of Resample in Algorithm 1, we perform a greedy resampling, i.e., we pick the best candidate within each group of size $g_{t}$ and replicate it $g_{t}$ times. Since we are using large enough $B$, this is justified and has similar effect as using smaller $B$ with moderate temperature, with the added advantage of utilizing less function calls to the reward network $r$. Thus, tuning the $B$ allows us to maintain balance without over-optimizing with respect to the reward.
For reproducibility, we provide detailed settings for each task, sampler, and search algorithm used in our experiments below.
C.1 Face Identity Experiments
For experiments with face identity as side information:
  • DPS: Box inpainting with box size $96$. All other parameters (downsampling rate, blur kernel, noise levels) are the same as the default DPS settings in the implementation. Search algorithms use $N=8$ particles and resampling base $B=16$ for both RFJ and Greedy search. Gradient guidance is applied with scale $0.5$.
  • DAPS: Box inpainting with box size $96$, super-resolution with downsampling rate $10\times$. Noise levels unchanged from original DAPS defaults. Search algorithms use $N=8$ particles and resampling base $B=4$, with gradient guidance scale $13$.
  • MPGD: Box inpainting with box size $64$, super-resolution with downsampling rate $6\times$, and Gaussian deblur with intensity of $5.0$. Search algorithms use $N=8$ particles and resampling base $B=8$. Gradient guidance scale is $0.5$ for box inpainting and $0.25$ for super-resolution and Gaussian deblur.
C.2 Text Side Information Experiments
When text descriptions were used as side information, we made the degradation more severe so that the information in $\mathbf{s}$ was not already present in the measurement $\mathbf{y}$. Otherwise, side information would not provide meaningful guidance. For example, if the input image is sharp enough to identify the type of animal, then explicitly stating it in $\mathbf{s}$ adds little value.
The settings for these tasks are:
  • Box inpainting: Box size $138$, noise level same as default.
  • Super-resolution: Downsampling rate $32\times$, noise level same as default.
  • Motion/Gaussian deblur: Kernel size $256$, intensity of Gaussian $5.0$, noise level $0.1$.
  • Nonlinear/Blind deblur: Kernel size unchanged, noise level $0.5$.
For all tasks in this setting, search algorithms use $N=4$ particles and resampling base $B=100$.
These hyperparameters ensure that our framework is evaluated under severe degradations (heavy downsampling, blurring, or noise), while search and guidance settings remain consistent across samplers and modalities.
C.3 MRI Experiments
We used the contrast-pairings among the files in the fastMRI dataset, provided by Atalık et al., (2025). We collect the data from the (fastMRI) source and preprocess to be compatible with the inputs in ContextMRI. Specifically, the setup used in the data is multi-coil MRI acquisition, which requires us to estimate the coil sensitivity maps, and then a complex reconstruction from them. ContextMRI takes complex values as inputs and denoises to produce a complex-valued 2D image. We computed NMI with $64$ bins at each step of the diffusion process to balance complexity with performance. We use the defaults parameters as in ContextMRI, except for the acceleration factor, $16$ and the center fraction (ACS), $0.02$. We use a pair of anatomy which two contrasts, which has more 30 slices. We consider the slices $15$ to $28$ as these are more challenging and report the results by using one as the side information for the other.
Appendix D Limitations
Our proposed search algorithms lack formal optimality guarantees for exploration–exploitation, and we do not claim they are theoretically optimal. We expect that stronger algorithms are possible, potentially improving both sample efficiency and robustness. A central reason is the absence of a general mathematical framework for designing optimal exploration–exploitation strategies in diffusion-based inverse problems with side information, an open problem we highlight. Practically, this means our methods rely on principled heuristics and tuned schedules (e.g., reward scaling, resampling/branching rates) whose compute allocation is not provably optimal, suggesting a clear direction for future work.
Appendix E Use of Large Language Models
Parts of this work were assisted by large language models (specifically GPT-5 from OpenAI). Their use was limited to improving clarity, grammar, and the presentation of experimental descriptions. All conceptual contributions, experimental design, analysis, and final decisions are solely the authors’ responsibility. The models were not used to generate new research ideas, design experiments, or make unverifiable scientific claims. Additionally, large language models were used to generate textual descriptions serving as side information for a specific set of experiments, as detailed in the main paper.